### v2

# TimeScope: Towards Task-Oriented Temporal Grounding In Long Videos

Abstract
Identifying key moments in long videos is essential for downstream understanding and reasoning tasks. In this paper, we introduce a new problem, Task-oriented Temporal Grounding (ToTG), which aims to localize time intervals containing the necessary information based on a task’s natural description. Along with the definition, we also present ToTG-Bench, a comprehensive benchmark for evaluating the performance on ToTG. ToTG is particularly challenging for traditional approaches due to their limited generalizability and difficulty in handling long videos. To address these challenges, we propose TimeScope, a novel framework built upon progressive reasoning. TimeScope first identifies a coarse-grained temporal scope in the long video that likely contains the key moments, and then refines this scope through fine-grained moment partitioning. Additionally, we curate a high-quality dataset, namely ToTG-Pile, to enhance TimeScope’s ability to perform progressive temporal grounding effectively. Extensive experiments demonstrate that TimeScope consistently outperforms both existing temporal-grounding methods and popular MLLMs across various settings, highlighting its effectiveness in addressing this new challenging problem.
1 Introduction
Temporal grounding is crucial for a wide range of applications Gao et al. (2017); Yuan et al. (2021), such as video question-answering and abnormality surveillance. It has been a long-standing research topic, and recent progresses on multi-modal large language models (MLLMs) have substantially advanced the study in this area Zeng et al. (2025); Ren et al. (2024); Wang et al. (2025c); Li et al. (2025e); Huang et al. (2024b); Guo et al. (2025); Shu et al. (2025); Yang et al. (2025c). However, most existing methods focus on temporal grounding based on explicit descriptions of the target. For example, a query like “identify the moment where a little boy is holding a basketball” clearly specifies the event to be localized, allowing models to perform grounding through straightforward visual–text alignment. In contrast, real-world applications often require a more implicit form of temporal grounding. In these scenarios, the model must localize key moments based on a task’s natural description, where the relevant information is not directly stated. For instance, given the task “why the boy looks happy when he comes home”, the key moment corresponds to the event “he receives a gift from his friend”. Although the importance for completing the task, its connection to the task description is indirect, making accurate grounding far more challenging.
In this paper, we introduce a new problem, called Task-oriented Temporal Grounding (ToTG), to formally conceptualize the aforementioned challenge. Specifically, given a task’s natural description, the goal of ToTG is to identify the time intervals within a video that contain the necessary information to solve the task. To facilitate research in this new area, we also create ToTG-Bench, a comprehensive benchmark designed to evaluate temporal grounding performance across a diverse set of real-world, long-video understanding scenarios. This benchmark provides a unified and challenging testbed for systematically comparing different approaches and accelerating progress in this emerging area.
The ToTG problem presents significant challenges for existing temporal grounding methods due to two key limitations. First, performing fine-grained grounding in long videos is inherently difficult, as models must sift through vast amounts of content filled with distracting and irrelevant information Li et al. (2025d). Second, current methods often struggle with generalizability, as they are typically trained to localize moments based on explicit event descriptions Anne Hendricks et al. (2017); Zala et al. (2023); Oncescu et al. (2021), rather than the implicit and diverse natural task descriptions encountered in real-world scenarios.
To address the above challenges, we propose TimeScope, a novel framework designed to tackle the ToTG problem through progressive reasoning. TimeScope operates in two stages to accurately localize crucial time intervals in long videos. First, it leverages abstracted video representations that capture comprehensive, high-level information about the entire video while intentionally sacrificing less important details. Using these abstract representations, TimeScope estimates a coarse-grained temporal scope, narrowing down the search to a subspace that are most likely to contain the needed information. Second, TimeScope re-loads detailed video representations for the selected scope and performs fine-grained partitioning to precisely localize the key moments within that region. This progressive process enables the model to effectively handle long videos while achieving high grounding accuracy. In addition to the core framework, we curate a new dataset, ToTG-Pile, specifically designed to optimize TimeScope’s performance on task-oriented temporal grounding. ToTG-Pile is sourced from diverse, real-world long-video datasets and annotated through a carefully designed pipeline which ensures both quality and diversity. TimeScope is trained on ToTG-Pile using a two-stage supervised fine-tuning strategy. In the first stage, we warm up the model’s temporal grounding capability by directly predicting target time intervals from the input video and task description. In the second stage, we establish the progressive reasoning capability by training the model to first estimate coarse-grained temporal scopes using abstracted video representations, and then refine its predictions to identify fine-grained time intervals based on detailed representations.
We conduct comprehensive experiments across a wide range of scenarios, evaluating TimeScope not only on ToTG-Bench but also on popular benchmarks for traditional temporal groundingCaba Heilbron et al. (2015); Sigurdsson et al. (2016); Cheng et al. (2025). The results demonstrate that TimeScope delivers significant improvements over existing approaches, including both dual-encoder-based model, specialized MLLMs for temporal grounding and advanced generic MLLMsBai et al. (2025); Yang et al. (2025a). Further analyses highlight the contributions of each component within our framework, validating its effectiveness in addressing this new and challenging problem. To facilitate future research, we will publicly release all resources, including model, dataset, benchmark, and source code.
2 Related work
2.1 Video Temporal Grounding
The traditional temporal grounding (TG) task requires models to localize a time interval in a video given a query that explicitly describes the target event. Early approaches are mainly dual-encoder-based, where video and language features are extracted using different pre-trained encoders (e.g., BERT Devlin et al. (2019), CLIP Radford et al. (2021), SigLip Zhai et al. (2023)), and then fused for time interval prediction Mu et al. (2024); Lei et al. (2021); Moon et al. (2023b; a); Gordeev et al. (2024). These models rely on simple visual-text alignment, which limits their generalization to out-of-domain or more complex queries.
More recently, researchers have explored applying MLLMs to temporal grounding Huang et al. (2024b); Ren et al. (2024); Zeng et al. (2025); Guo et al. (2024); Wang et al. (2025b); Li et al. (2025f). For instance, TimeChat Ren et al. (2024) introduces a time-aware frame encoder that binds visual tokens with their corresponding timestamps at the frame level for temporal grounding. Similarly, TimeSuite Zeng et al. (2025) proposes temporal-adaptive position encoding to strengthen temporal awareness in video representations. Trace Guo et al. (2024) designs a specialized encoder and head for timestamp input, while Time-R1 Wang et al. (2025b) employs a reasoning-guided post-training framework with reinforcement learning and verifiable rewards to improve grounding accuracy. In addition to these specialized MLLMs, recent generic MLLMs (e.g., Qwen2.5-VL Bai et al. (2025), Keye-VL-1.5 Yang et al. (2025b)) have also demonstrated certain capabilities for temporal grounding.
Despite these advances, existing methods are still limited in handling complex queries that require not just locating explicit events but identifying intervals that support completing a task in long video. In particular, most MLLM-based approaches underutilize the generalization potential of MLLMs. Motivated by these limitations, we introduce the new problem of task-oriented temporal grounding, along with a benchmark, a dataset, and a dedicated framework to address it.
2.2 Long Video Understanding
The field of long video understanding (LVU) has developed rapidly in recent years, with many powerful MLLMs emerging, such as VideoChatFlash Li et al. (2025b), Video-XL-2 Qin et al. (2025a), Eagle2.5 Chen et al. (2025), and InternVL3 Zhu et al. (2025). These models demonstrate strong general video understanding capabilities and serve as versatile backbones for various video tasks, including temporal grounding.
However, precisely localizing or perceiving fine-grained details within second-level intervals remains a major challenge for current LVU MLLMs. To address this, some works introduce additional modules to identify key frames or video segments based on task-oriented queries Wang et al. (2025a); Huang et al. (2025); Yu et al. (2024); Qin et al. (2025b). These modules are typically similarity-based and thus lack deeper semantic understanding of the video content, limiting their compatibility with diverse downstream tasks in long video scenarios.
In contrast, we take a different approach. We post-train LVU MLLMs on our diverse and high-quality task-oriented grounding dataset, and further implement TimeScope, a novel framework designed for progressive task-oriented grounding. This enables the model to efficiently and accurately localize critical time intervals in long videos for a wide range of tasks.
3 Method
In this section, we introduce our new proposed problem, Task-Oriented Temporal Grounding (ToTG) along with its associated benchmark ToTG-Bench, grounding framework TimeScope and training dataset ToTG-Pile in detail. In the following, Section 3.1 formulates the ToTG, which is a more challenging and valuable task compared to traditional temporal grounding. Section 3.2 presents ToTG-Bench, a comprehensive benchmark designed to evaluate model in localizing time intervals based on task description across various long video scenarios. Section 3.3 details our novel framework TimeScope, which aims to address the challenges of TOTG via progressive reasoning. Section 3.4 introduces the dataset TOTG-Piles and its carefully designed pipeline.
3.1 Problem Formulation
We formally define the Task-oriented Temporal Grounding (ToTG) problem in this section. Given an untrimmed video and a task-oriented query , the goal is to localize a temporal interval within that contains the information necessary to accomplish the task described by . In conventional temporal grounding, the query is usually an event description, including explicit information of the grounding target. This enables the model to locate the corresponding time interval via relatively straightforward visual-text alignment. In contrast, the query in ToTG supplements a higher-level task instruction about the target on this basis. Therefore, the model must locate the significant time interval by comprehensively understanding both visual and textual content, rather than relying solely on surface-level alignment.
3.2 Benchmarks
To facilitate the evaluation of ToTG, we introduce ToTG-Bench, a comprehensive and challenging benchmark. Unlike conventional temporal grounding benchmarks Sigurdsson et al. (2016); Caba Heilbron et al. (2015), ToTG-Bench features queries spanning 12 distinct task types. Its videos are drawn from diverse long-video understanding scenarios and cover durations ranging from a few seconds to nearly one hour. We construct the benchmark through a combination of an effective data filtering pipeline and careful manual annotation, ensuring accurate temporal intervals for each test sample. All of these make ToTG-Bench a high-quality, diverse, and challenging testbed for advancing research in ToTG.
Benchmark Construction We first collect samples from four long-video understanding benchmarks (MLVU, Video-MME, LongVideoBench, and V-STaR) as the candidate data for ToTG-Bench construction. The questions in these samples are used as task-oriented queries for grounding. However, not all data are suitable for ToTG, so we design the following filtering pipeline to obtain qualified candidates: 1) Task type filtering. We exclude samples whose task types are not compatible with ToTG (e.g., multilingual, summarization, event-ordering tasks and etc.). 2) Uniqueness filtering. We retain only samples where the query corresponds to a unique single grounding target. Specifically, we divide each long video into segments and evaluate them using a temporal grounding model followed by manual verification. Samples with exactly only one valid interval contained target are retained. 3) Information validation. We further verify that the preserved intervals contain sufficient information to answer the query. We use a advanced MLLM to generate answers based on sampled frames from the predicted intervals, and discard samples for which the answers fail. This pipeline yields a set of qualified candidate data with unique grounding targets, ensuring both annotation quality and fairness in subsequent evaluation. Furthermore, we balance the distribution of candidate data across query task types, video categories, video durations, and temporal positions of the target intervals (We refer to this as Query Center) so that each time interval within videos has a roughly equal chance of containing the target. Finally, we carefully manual annotate time interval to ensure the quality of each sample.
Benchmark Statistics As illustrated in Figure 2, ToTG-Bench demonstrates significant diversity across task types, video categories, video durations and position of target interval. Specifically, it includes 12 task types (e.g., action reasoning, OCR perception temporal reasoning and etc.), covering a broad range of video understanding tasks 111We follow the definition of task type and video category from Video-MME. It further comprises 35 video categories, covering a wide range of real-world video domains. The video durations in ToTG-Bench range from a few minutes to up to one hour. Moreover, the target intervals are balanced to be uniformly distributed along the video timeline, ensuring unbiased evaluation.
3.3 TimeScope
ToTG requires models to identify the necessary information needed to accomplish diverse tasks in long videos. However, existing temporal grounding models are often distracted by abundant irrelevant content in long contexts, which prevents them from accurately grounding complex task-oriented queries. To overcome this, we design TimeScope, a progressive reasoning framework that adapts MLLM backbones for task-oriented temporal grounding. Instead of directly predicting over the full long video at once, TimeScope performs reasoning in two complementary stages, progressively refining from holistic context to precise details (Figure 1).
Stage 1 (Coarse reasoning). TimeScope first compresses the cached Key-Value (KV) states of visual tokens (detonated as ) obtained during pre-filling. Through average pooling, the fine-grained are distilled into , which serves as a compact representation of the global video context. On this compact KV space, the model performs a holistic reasoning step to hypothesize a coarse temporal window that is likely to contain the target.
Stage 2 (Fine reasoning). Given the temporal window , TimeScope selectively reloads only the fine-grained from cache for the corresponding frames, while discarding irrelevant context outside . This selective refinement enables the model to reason over detailed visual cues within the localized window, and to output a precise grounding interval.
By framing temporal grounding on long frames as a progressive reasoning process, TimeScope achieves both superior efficiency and accuracy in task-oriented long video temporal grounding.
3.4 Datasets
Existing temporal grounding datasets are mostly limited to short videos with explicit event queries, which makes models trained on them poorly suited for ToTG task. To bridge this gap, we curate a new large-scale dataset, ToTG-Pile, specifically designed for task-oriented temporal grounding. ToTG-Pile is diverse and comprehensive, incorporating both traditional temporal grounding data and newly constructed task-oriented data.
Traditional temporal grounding data. Existing datasets for temporal grounding are limited in terms of video duration. To address this, we recaption existing short-video datasets using Qwen2.5-VL to generate concise descriptions. We then concatenate these short videos into longer videos, and the concise descriptions are used to construct explicit temporal grounding queries. Through this pipeline, we obtain a training set of 85k long-video temporal grounding samples, with an average video duration of about 500 seconds (approximately 8 minutes).
Task-oriented temporal grounding data. For ToTG data, we collect raw samples from NextQA Xiao et al. (2021), STAR Wu et al. (2024), and CLEVRER Yi et al. (2020). As illustrated in Figure 1(b), the video and answer of each sample are fed into multiple temporal grounding models (refer to Video Expert) to generate candidate time intervals containing the answer. A sample is retained only if the Intersection-over-Union (IoU) between intervals predicted by different models exceeds 0.5, and is further evaluated by another MLLM to ensure it contains sufficient information to answer the question. The resulting qualified QA pairs with timestamps are then concatenated into longer videos, yielding training data tailored for task-oriented grounding.
Overall, ToTG-Pile unifies traditional grounding data with newly constructed task-oriented data, ensuring diversity across tasks, durations, and video domains. ToTG-Pile lays the foundation for developing excellent temporal grounding models in ToTG task.
4 Experiment
4.1 Implementation details
We adopt VideoXL-2 as our backbone for two reasons: first, it can ingest very long video sequences, which makes it straightforward to build long-video temporal-understanding methods; second, its internal design interleaves timestamp tokens, giving the model a strong built-in sense of time. TimeScope is trained on ToTG-Pile with a two-stage supervised fine-tuning schedule. In Stage 1 we use the temporal-grounding splits of ToTG-Pile and ask the model to predict the target time interval directly from the raw video and the task description, thereby bootstrapping its basic localization ability. In Stage 2 we apply heavy temporal augmentations—random cropping, shifting, and scaling of the time span—to the training videos, forcing the model to first estimate a coarse temporal window from an abstract video representation and then refine that window into a fine-grained interval using the detailed representation. Throughout training we sample video frames at 1 fps, capping at 300 frames maximum.
| Method | Charades-STA | ActivityNet | ||||
| R1@0.5 | R1@0.7 | mIoU | R1@0.5 | R1@0.7 | mIoU | |
| Open-source VLP Method | ||||||
| 2D-TAN Zhang et al. (2020) | 45.8 | 27.9 | – | 60.4 | 43.4 | – |
| UniVTG Lin et al. (2023) | 60.2 | 38.6 | – | 56.1 | 43.4 | – |
| SSRN Zhu et al. (2022) | 65.5 | 42.6 | – | – | 54.5 | – |
| SnAG Mu et al. (2024) | 64.6 | 46.2 | – | – | 48.6 | – |
| EaTR Jang et al. (2023) | 68.4 | 44.9 | – | – | 58.2 | – |
| Open-source MLLMs Method | ||||||
| TimeChat Ren et al. (2024) | 32.2 | 13.4 | 32.2 | 36.2 | 20.2 | 21.8 |
| VTimeLLM Huang et al. (2024a) | 27.5 | 11.4 | 31.2 | 44.0 | 27.8 | 30.4 |
| VideoChat-Flash Li et al. (2024) | 53.1 | 27.6 | – | – | – | – |
| TRACE Guo et al. (2024) | 61.7 | 41.4 | 41.4 | 37.7 | 24.0 | 39.0 |
| HawkEye Wang et al. (2024) | 58.3 | 28.8 | – | 55.9 | 34.7 | – |
| TimeSuite Zeng et al. (2025) | 67.1 | 43.0 | – | – | – | – |
| Time-R1 Wang et al. (2025b) | 72.2 | 50.1 | – | 58.6 | 39.0 | – |
| DeepVideo-R1-7B Park et al. (2025) | 71.7 | 50.6 | 61.2 | 33.9 | 18.0 | 36.9 |
| VideoChat-R1-7B Li et al. (2025c) | 71.7 | 50.2 | 60.8 | 33.4 | 17.7 | 36.6 |
| TimeZero-7B Wang et al. (2025b) | 60.8 | 35.3 | 58.1 | 39.0 | 21.4 | 40.5 |
| Temporal-RLT-7BLi et al. (2025a) | 67.9 | 44.1 | 57.0 | 38.4 | 20.2 | 39.0 |
| TimeScope-7B | 78.9 | 64.0 | 57.5 | 69.6 | 59.0 | 48.0 |
4.2 Benchmarks and Metrics
For evaluation, we benchmark our method across the following four categories: (i) short-video temporal grounding, including Charades-STA [46] and ActivityNet-Captions; (ii) long-video temporal grounding, including videos longer than 300s from V-STaR and Vid-Chapters-7M; (iii) Video Question Answering (VideoQA), including three general long-video QA benchmarks: CG-Bench, MLVU, and LongVideoBench ; (iv) long video task grounding, including our proposed ToTG-bench. The statistics of the evaluation benchmarks used are listed in Table 5.
Evaluation metrics. For temporal grounding tasks and task grounding tasks, we adopt Recall@1 (R@1) at multiple temporal intersection-over-union (IoU) thresholds and mean IoU (mIoU) as evaluation metrics. Specifically, for temporal grounding benchmarks, we use IoU thresholds of 0.3, 0.5, and 0.7. For general VideoQA tasks, we report standard accuracy metrics.
| Model | R1@0.5 | R1@0.7 |
| VTimeLLM-7B Huang et al. (2024a) | 4.1 | 1.6 |
| CLIP Radford et al. (2021) | 5.2 | 2.3 |
| M-DETR Kamath et al. (2021) | 27.3 | 17.6 |
| ReVisionLLM Hannan et al. (2024) | 27.4 | 21.8 |
| Qwen2.5-VL-7B Bai et al. (2025) | 0.3 | 0.1 |
| VITAL-7B Zhang et al. (2025) | 25.8 | 19.5 |
| TimeScope-7B | 24.3 | 22.1 |
| Model | R1@0.5 | R1@0.7 |
| Qwen2.5-VL-7B | 0.0 | 0.0 |
| UniTime | 0.629 | 0.629 |
| Keye-1.5-VL-8B | 0.639 | 0.491 |
| TimeScope-7B | 0.909 | 0.909 |
| Method | CG-Bench | MLVU | LongVideoBench |
| Acc. | Acc. | Acc. | |
| Uniform Sample | 33.87 | 60.53 | 54.82 |
| UniVTG Lin et al. (2023) | 34.87 | 62.56 | 54.67 |
| VTimeLLM Huang et al. (2024a) | 34.60 | 59.52 | 54.30 |
| TimeSuite Zeng et al. (2025) | 32.47 | 58.51 | 53.25 |
| UniTime-Full Li et al. (2025f) | 40.30 | 66.50 | 56.47 |
| TimeScope-7B | 39.53 | 70.14 | 59.18 |
| Method | S(R1@0.5) | S(R1@0.7) | M(R1@0.5) | M(R1@0.7) | L(R1@0.5) | L(R1@0.7) |
| no option | ||||||
| Qwen-2.5VL-7B | 0.458 | 0.400 | 0.111 | 0.091 | 0.030 | 0.030 |
| ARC-Hunyuan-Video-7B | 0.523 | 0.484 | 0.36 | 0.260 | 0.175 | 0.162 |
| Keye-VL-1.5-8B | 0.521 | 0.333 | 0.313 | 0.261 | 0.221 | 0.124 |
| UniTime | 0.524 | 0.427 | 0.296 | 0.232 | 0.242 | 0.226 |
| TimeScope-7B | 0.574 | 0.436 | 0.47 | 0.44 | 0.435 | 0.405 |
| has option | ||||||
| Qwen-2.5VL-7B | 0.475 | 0.413 | 0.111 | 0.091 | 0.030 | 0.030 |
| ARC-Hunyuan-Video-7B | 0.158 | 0.148 | 0.111 | 0.090 | 0.124 | 0.092 |
| Keye-1.5-VL-8B | 0.527 | 0.363 | 0.514 | 0.471 | 0.261 | 0.185 |
| UniTime | 0.557 | 0.441 | 0.317 | 0.274 | 0.287 | 0.256 |
| TimeScope-7B | 0.683 | 0.515 | 0.52 | 0.44 | 0.467 | 0.426 |
4.3 SOTA Performance on Traditional Temporal Grounding
We conduct a comprehensive comparison of TimeScope against dual-encoder-based methods, MLLM-based methods, and video understanding models on both short- and long-video temporal grounding benchmarks. As shown in Table 1, TimeScope achieves state-of-the-art (SOTA) performance across all benchmarks. For instance, on Charades-STA, TimeScope reached an R1@0.7 score of 64.0, surpassing VideoChat-Flash (27.6), TimeSuite (43.0), and Time-R1 (50.1). On ActivityNet, it achieved an R1@0.7 score of 59, outperforming HawkEye (34.7) and Time-R1 (39.0).
Particularly noteworthy is that, compared to most other models, TimeScope maintains a smaller gap between R@1@0.5 and R@1@0.7. This indicates that our model is capable of more precise temporal localization. Furthermore, as shown in Table 3, TimeScope achieved an R1@0.7 score of 90.9 on V-STaR, where video duration reaches 300 seconds, significantly surpassing UniTime’s 62.9 and Keye-1.5-VL’s 49.1, demonstrating its superior performance and strong capability in handling long-video grounding tasks.
4.4 SOTA Performance on Task-oriented temporal grounding
Besides traditional temporal grounding, we also evaluate TimeScope on ToTG-Bench, comparing it with MLLM-based models and generic MLLMs. As shown in Table LABEL:lvtgbench, where “has option” indicates that the options are fed into the model along with certain prompt information, TimeScope achieves outstanding and robust performance across different video durations. Specifically, while TimeScope and ARC-Hunyuan-Video-7B Ge et al. (2025) achieve comparable performance on short videos (within a 5-point difference), TimeScope shows clear superiority on medium and long videos, where it outperforms other models by 10 to 20 points. This indicates that our approach can effectively handle videos of various durations, especially long videos, and demonstrates excellent grounding ability for task-oriented queries.
| Model | Early | Middle | Late | Std Dev () |
| Qwen-2.5-VL Bai et al. (2025) | 20.0 | 13.4 | 4.1 | 8.0 |
| Keye-1.5-VL Yang et al. (2025a) | 50.0 | 33.3 | 31.1 | 10.4 |
| UniTime-Full Yang et al. (2025a) | 50.9 | 28.1 | 44.0 | 11.6 |
| TimeScope-7B | 39.3 | 44.8 | 46.0 | 3.5 |
4.5 Improvement for Long-Video QA
As discussed in Section 3.1, the strong performance of TimeScope on the ToTG task suggests that it can help MLLMs capture the critical information in long videos required for question answering. To validate this point, we first use TimeScope and other temporal grounding models to localize time interval. The frames within the predicted intervals are then fed into Qwen2.5-VL-7B for answering. Their results are compared against the default setting, where frames are uniformly sampled without temporal grounding. We evaluate this experiment on three long-video understanding benchmarks: CG-Bench, MLVU, and LongVideoBench. As shown in Table 4(a), TimeScope consistently brings significant improvements across all benchmarks, while other temporal grounding models fail to outperform the default uniform-sampling baseline.
| Model | R1@0.3 | R1@0.5 | R1@0.7 |
| TimeScopeScope | |||
| Zeroshot | 89.7 | 85.2 | 73.8 |
| TimeScopeScope | |||
| Multi-stage | 92.1 | 90.9 | 90.9 |
| Method | Prefill | Decode | Cost |
| zero shot | |||
| Stage1 | 1400ms | 672ms | 2072ms |
| Multi-stage | |||
| Stage1 | 1400ms | 441ms | |
| Stage2 | 158ms | 357ms | 2356ms |
4.6 Ablation Studies
Effectiveness of Progressive Reasoning. To assess the effectiveness and necessity of progressive reasoning, we conduct an ablation study comparing two settings: progressive reasoning and single-step reasoning. The study is performed on videos longer than 300 seconds from the V-STaR benchmark, with results reported in Table 4(a). We observe that progressive reasoning achieves substantial improvements over the single-step baseline. In particular, for longer videos, progressively narrowing the search space proves both effective and necessary for accurate temporal grounding.
Efficiency of Progressive Reasoning. In addition to effectiveness, we also evaluate the computational efficiency of TimeScope. As shown in Table 4(b), progressive reasoning introduces only marginal additional overhead compared to single-step reasoning, while still delivering substantial accuracy gains. This efficiency largely benefits from our design of adjusting KV sparsity during progressive reasoning. Overall, TimeScope strikes a cost-effective balance between accuracy and computational cost, making it practical for long-video temporal grounding.
Robustness against Time Bias. After analyzing the evaluation results on ToTG-Bench, we find that the Query Center (the position of the target time interval) introduces a significant bias that strongly influences the performance of temporal grounding models. As shown in Table 6, Qwen2.5-VL and Keye-1.5-VL tend to perform better when the target time interval appears at the beginning of the video, but their performance drops substantially when the target lies in the middle or towards the end. In contrast, our TimeScope model maintains robust performance across different target interval positions, demonstrating its resilience against this bias.
5 Conclusion
In this work, we define a new task—Task-Oriented Temporal Grounding (ToTG)—and formally conceptualize the aforementioned challenges. To foster research in this emerging area, we introduce ToTG-Bench, a benchmark designed to evaluate temporal grounding performance on diverse, real-world, long-form video-understanding scenarios. To tackle these challenges, we propose TimeScope, a novel framework that solves ToTG through step-by-step reasoning. To strengthen TimeScope, we release ToTG-Pile, a dataset expressly engineered to optimize MLLMs for task-oriented temporal grounding. Harvested from diverse real-world long-video corpora and annotated via a carefully engineered pipeline, ToTG-Pile provides large-scale, high-quality training data. Extensive experiments across a wide spectrum of settings show that TimeScope achieves substantial improvements over existing methods on both traditional benchmarks and ToTG-Bench. We hope this work will stimulate future research on Task-Oriented Temporal Grounding and propel MLLMs toward deeper temporal understanding of video.
Appendix A Use of LLMs
In the process of writing this paper, we utilized a Large Language Model (LLM) solely as a tool for polishing the language and enhancing the readability of the manuscript. The LLM was employed to refine the wording, grammar, and coherence of the text, ensuring that the content is clear and well-structured. However, it is important to note that the LLM did not play any significant role in the conception of research ideas or the formulation of the paper’s content. All research ideas, methodologies, and conclusions presented in this paper are the original work of the authors. We understand that we bear full responsibility for the contents of this paper, including any text that has been processed by the LLM. We have carefully reviewed and verified all content to ensure its accuracy and originality. Any contributions made by the LLM were limited to language enhancement and did not involve any form of plagiarism or scientific misconduct.
Appendix B Benchmark Details
To evaluate the performance of our model on temporal grounding and VideoQA tasks, we employ the Intersection-over-Union (IoU) metric and its variants. These metrics quantify the alignment between the predicted temporal window and the ground truth temporal window . The formal definitions are as follows:
B.1 IoU and mIoU
The fundamental IoU metric is defined as:
The mean Intersection-over-Union (mIoU) is calculated as the average IoU across all test samples:
where is the total number of test samples and is the IoU value for the -th sample.
We also evaluate performance using IoU thresholds, which measure the percentage of predictions exceeding specific IoU values:
IoU@0.3: Percentage of predictions with IoU 0.3
IoU@0.5: Percentage of predictions with IoU 0.5
IoU@0.7: Percentage of predictions with IoU 0.7
Here, denotes the duration of the overlapping region between the predicted window and the ground truth window . represents the duration of their union, while and denote the durations of the predicted and ground truth windows, respectively. The indicator function equals 1 when the condition is true and 0 otherwise.
IoU measures the overall alignment between and , providing a balanced assessment of both precision and recall by considering the overlap relative to their union. The threshold-based metrics (IoU@0.3, IoU@0.5, IoU@0.7) evaluate the model’s ability to produce high-quality predictions meeting different precision standards, while mIoU provides an overall average performance measure across all samples.
B.2 Query Center Robustness
To measure the effectiveness of balanced query centers in the ToTG-benchmark, we evaluated the sensitivity of state-of-the-art (SOTA) models to query centers. As shown in the figure, experiments indicate that most models perform better on test data with query centers positioned earlier rather than later or in the middle. We measured the difference between the best and worst performance of various models when the query center varies. The results show that Timescope maintains its effectiveness regardless of the position of the query center (for instance, the gap is only 35% at iou@0.3, which is significantly better than Qwen2.5vl’s 78% and 23.8% higher than UniTime). This further proves the effectiveness of Timescope’s training.
| Query Center | 0-0.1 | 0.1-0.2 | 0.2-0.3 | 0.3-0.4 | 0.4-0.5 | 0.5-0.6 | 0.6-0.7 | 0.7-0.8 | 0.8-0.9 | 0.9-1 | diff | |
| Qwen2.5VL | R@0.3 | 0.417 | 0.368 | 0.615 | 0.222 | 0.250 | 0.158 | 0.278 | 0.235 | 0.133 | 0.154 | 78% |
| R@0.5 | 0.250 | 0.158 | 0.462 | 0.185 | 0.062 | 0.158 | 0.111 | 0.011 | 0.066 | 0.000 | 97% | |
| R@0.7 | 0.208 | 0.158 | 0.308 | 0.074 | 0.000 | 0.053 | 0.055 | 0.000 | 0.000 | 0.000 | 100% | |
| Keye-1.5-VL | R@0.3 | 0.615 | 0.652 | 0.666 | 0.444 | 0.375 | 0.353 | 0.556 | 0.600 | 0.235 | 0.461 | 64% |
| R@0.5 | 0.538 | 0.521 | 0.400 | 0.379 | 0.186 | 0.294 | 0.444 | 0.466 | 0.176 | 0.307 | 67.2% | |
| R@0.7 | 0.462 | 0.260 | 0.266 | 0.259 | 0.125 | 0.176 | 0.222 | 0.333 | 0.117 | 0.231 | 74.6% | |
| Unitime | R@0.3 | 0.669 | 0.465 | 0.633 | 0.364 | 0.371 | 0.216 | 0.550 | 0.567 | 0.371 | 0.567 | 67.7% |
| R@0.5 | 0.592 | 0.335 | 0.633 | 0.221 | 0.294 | 0.216 | 0.500 | 0.566 | 0.311 | 0.433 | 65.0% | |
| R@0.7 | 0.476 | 0.335 | 0.367 | 0.186 | 0.176 | 0.163 | 0.500 | 0.455 | 0.194 | 0.433 | 67.4% | |
| Timescope | R@0.3 | 0.393 | 0.417 | 0.606 | 0.464 | 0.421 | 0.470 | 0.555 | 0.667 | 0.470 | 0.467 | 35.1% |
| R@0.5 | 0.393 | 0.417 | 0.606 | 0.357 | 0.368 | 0.353 | 0.505 | 0.556 | 0.412 | 0.400 | 41.7% | |
| R@0.7 | 0.357 | 0.375 | 0.533 | 0.286 | 0.263 | 0.294 | 0.400 | 0.500 | 0.353 | 0.267 | 50.6% | |
Appendix C Model Arch and Training detail
Timescope initializes with the base model of VideoXL2 because it can handle very long sequences, facilitating the construction of long video Task-oriented Temporal Grounding tasks. The overall architecture is shown in the figure above. The architecture of Timescope consists of four components: a visual encoder, Dynamic Token Synthesis (DTS), an MLP projector, and a Large Language Model (LLM). For the visual encoder, we use SigLIP to encode visual inputs into dense visual features. The DTS module, located after the visual encoder, is constructed by combining spatial-temporal attention blocks and 3D convolutional layers. It aims to process visual features extracted from four consecutive visual inputs as a single group. This design has been validated in Video-XL-Pro. Next, the architecture applies average pooling to adjacent features to further compress the representation. Then, a two-layer MLP projector processes these pooled features, projecting them into the embedding space of the LLM. For our LLM, we adopt Qwen2.5-7B.
Furthermore, in Timescope, to enhance the model’s temporal understanding capability, we prepend explicit timestamp tokens (e.g., Time: 4.0 Second) for every four consecutive frames group across the entire frames sequence. This direct timing information substantially improves the model’s temporal awareness. During training, we randomly perform shift operations (e.g., 4s to 1004 seconds) and cut operations (e.g., cutting out 10-20 seconds of a video as a single data point) on the explicit timestamp tokens of the input video as a whole. This training technique greatly enhances Timescope’s multi-stage understanding capability and robustness to query positions, as shown in Table B.2.
Appendix D Experimental Settings & Additional Results
We elaborate on the training and inference details of Timescope. The reported hyperparameters cover Stage 1, and 2, as specified in Table 1. For the inference details, we emphasize the particular context length for different benchmarks, as shown in Table 2.
| Hyperparameter | Stage 1 | Stage 2 |
| Overall batch size | 64 | 64 |
| Learning rate | 1e-5 | 1e-5 |
| LR Scheduler | Cosine decay | Cosine decay |
| DeepSpeed ZeRO Stage | ZeRO-2-offload | ZeRO-2-offload |
| Optimizer | Adam | Adam |
| Warmup ratio | 0.3 | 0.3 |
| Epoch | 1 | 1 |
| Weight decay | 0 | 0 |
| Precision | bf16 | bf16 |
| Dataset | Context Length |
| Charades-STA | 1fps |
| Activity-Net | 1fps |
| Vid-Chapters | 1fps(800 frames) |
| CG-Bench | 1fps(800 frames) |
| MLVU | 1fps(800 frames) |
| LongVideoBench | 1fps(800 frames) |
| V-STaR | 1fps(800 frames) |
Appendix E VideoQA Task Details
E.1 VideoQA with Temporal Grounding
We use Qwen2-VL-7B as the VideoQA model for answer generation. By default, it processes long videos by uniformly sampling 32 frames. However, this sampling strategy may lead to the omission of critical information. To investigate whether temporal grounding models can compensate for this issue, we adopt the following procedure. First, we use different video temporal grounding models to localize the relevant segments for each question. Then, we crop the localized video intervals and input them into Qwen2-VL-7B. Specifically, for cropped video segments shorter than 32 seconds, we extend their duration from the center to 32 seconds. Within each interval, we again uniformly sample 32 frames for answer generation.
E.2 Prompt template for VideoQA
We use the same prompt template for all multiple-choice VideoQA benchmarks:
System: You are a helpful assistant. User: <video> Question: <question> Options: (A) <Option_A> (B) <Option_B> (C) <Option_C> (D) <Option_D> Please only give the best option. Best Option: Assistant:
Appendix F Limitations & Future Work
Although Timescope demonstrates exceptional performance on various video temporal grounding and video QA benchmarks, it still has several limitations that warrant further exploration: (i) Timescope is currently constrained to temporal grounding tasks (including traditional temporal grounding tasks and Task-Oriented Temporal Grounding tasks). To enable broader applications in MLLMs, it requires more diverse training data with dense temporal annotations. Incorporating such data into the pretraining process of MLLMs could unlock their potential for handling more temporally complex tasks, such as dense video captioning. (ii) Although Timescope enhances MLLMs with temporal grounding capabilities, relying solely on temporal grounding data limits their reasoning and question-answering abilities. The ultimate objective is to develop MLLMs that seamlessly integrate localization, reasoning, and question-answering into a unified framework.
Appendix G Broader Impacts Statement
Our research introduces a new problem, called Task-oriented Temporal Grounding (ToTG), to formally conceptualize the aforementioned challenge. We have created ToTG-Bench, aimed at evaluating temporal grounding performance in diverse real-world long video understanding scenarios and accelerating progress in this emerging field. This facilitates advancing the complexity of temporal understanding in videos and accelerates the development of models that integrate thinking and traditional temporal grounding capabilities. We hope that ultimately, the two can be integrated, unifying problem thinking and temporal grounding. We have also developed a more accurate and efficient temporal grounding framework, Timescope, to advance the field of long video temporal understanding. This could benefit a wide range of downstream applications, such as anomaly detection, security monitoring, etc. We believe that ToTG and Timescope can advance the development of video temporal understanding.
Appendix H Qualitative Results
In Figure 2–4, we present the qualitative results of Timescope on ToTG-bench and V-STaR. Timescope demonstrates accurate understanding of questions and the ability to provide temporal grounding. In V-STaR, we show the results when two-stage reasoning is applied, and it can be seen that Timescope exhibits robust performance with good coarse-grained segment retrieval and fine-grained temporal grounding capabilities.

### v3

@@
 Abstract
-Identifying key moments in long videos is essential for downstream understanding and reasoning tasks.
-In this paper, we introduce a new problem, Task-oriented Temporal Grounding (ToTG), which aims to localize time intervals containing the necessary information based on a task’s natural description.
-Along with the definition, we also present ToTG-Bench, a comprehensive benchmark for evaluating the performance on ToTG.
-ToTG is particularly challenging for traditional approaches due to their limited generalizability and difficulty in handling long videos.
-To address these challenges, we propose TimeScope, a novel framework built upon progressive reasoning.
-TimeScope first identifies a coarse-grained temporal scope in the long video that likely contains the key moments, and then refines this scope through fine-grained moment partitioning.
-Additionally, we curate a high-quality dataset, namely ToTG-Pile, to enhance TimeScope’s ability to perform progressive temporal grounding effectively.
-Extensive experiments demonstrate that TimeScope consistently outperforms both existing temporal-grounding methods and popular MLLMs across various settings, highlighting its effectiveness in addressing this new challenging problem.
+Identifying key temporal intervals within long videos, known as temporal grounding (TG), is important to video understanding and reasoning tasks.
+In this paper, we introduce a new form of the temporal grounding problem, Task-oriented Temporal Grounding (ToTG), which is driven by the requirements of downstream tasks rather than explicit time-interval descriptions.
+For example, a ToTG input may be “explain why the man in the video is sent to the hospital,” whereas traditional TG would take an explicit temporal description such as “the moments when the man is tripped by a stone and falls to the ground.” This new ToTG formulation presents significant challenges for existing TG methods, as it requires jointly performing deep task comprehension and fine-grained temporal localization within long videos.
+To address these challenges, we conduct a systematic set of studies.
+First, we construct a new benchmark ToTG-Bench, which comprehensively evaluates ToTG performance across diverse settings.
+Second, we introduce a new temporal-ground method TimeScope, which performs coarse-to-fine localization through a progressive reasoning process.
+Leveraging extensive supervised fine-tuning with carefully curated chain-of-thought (CoT) data from a variety of scenarios, TimeScope generalizes effectively across tasks and domains.
+Our evaluation demonstrates TimeScope’s empirical advantages over existing baselines from three perspectives: (1) substantial improvements in grounding precision, (2) significant benefits to downstream tasks, and (3) strong generalizability across different scenarios.
+All models, datasets, and source code will be fully open-sourced to support future research in this area.
 1 Introduction
-Temporal grounding is crucial for a wide range of applications Gao et al.
-(2017); Yuan et al.
-(2021), such as video question-answering and abnormality surveillance.
-It has been a long-standing research topic, and recent progresses on multi-modal large language models (MLLMs) have substantially advanced the study in this area Zeng et al.
-(2025); Ren et al.
-(2024); Wang et al.
-(2025c); Li et al.
-(2025e); Huang et al.
-(2024b); Guo et al.
-(2025); Shu et al.
-(2025); Yang et al.
-(2025c).
-However, most existing methods focus on temporal grounding based on explicit descriptions of the target.
-For example, a query like “identify the moment where a little boy is holding a basketball” clearly specifies the event to be localized, allowing models to perform grounding through straightforward visual–text alignment.
-In contrast, real-world applications often require a more implicit form of temporal grounding.
-In these scenarios, the model must localize key moments based on a task’s natural description, where the relevant information is not directly stated.
-For instance, given the task “why the boy looks happy when he comes home”, the key moment corresponds to the event “he receives a gift from his friend”.
-Although the importance for completing the task, its connection to the task description is indirect, making accurate grounding far more challenging.
-In this paper, we introduce a new problem, called Task-oriented Temporal Grounding (ToTG), to formally conceptualize the aforementioned challenge.
-Specifically, given a task’s natural description, the goal of ToTG is to identify the time intervals within a video that contain the necessary information to solve the task.
-To facilitate research in this new area, we also create ToTG-Bench, a comprehensive benchmark designed to evaluate temporal grounding performance across a diverse set of real-world, long-video understanding scenarios.
-This benchmark provides a unified and challenging testbed for systematically comparing different approaches and accelerating progress in this emerging area.
-The ToTG problem presents significant challenges for existing temporal grounding methods due to two key limitations.
-First, performing fine-grained grounding in long videos is inherently difficult, as models must sift through vast amounts of content filled with distracting and irrelevant information Li et al.
-(2025d).
-Second, current methods often struggle with generalizability, as they are typically trained to localize moments based on explicit event descriptions Anne Hendricks et al.
-(2017); Zala et al.
-(2023); Oncescu et al.
-(2021), rather than the implicit and diverse natural task descriptions encountered in real-world scenarios.
-To address the above challenges, we propose TimeScope, a novel framework designed to tackle the ToTG problem through progressive reasoning.
-TimeScope operates in two stages to accurately localize crucial time intervals in long videos.
-First, it leverages abstracted video representations that capture comprehensive, high-level information about the entire video while intentionally sacrificing less important details.
-Using these abstract representations, TimeScope estimates a coarse-grained temporal scope, narrowing down the search to a subspace that are most likely to contain the needed information.
-Second, TimeScope re-loads detailed video representations for the selected scope and performs fine-grained partitioning to precisely localize the key moments within that region.
-This progressive process enables the model to effectively handle long videos while achieving high grounding accuracy.
-In addition to the core framework, we curate a new dataset, ToTG-Pile, specifically designed to optimize TimeScope’s performance on task-oriented temporal grounding.
-ToTG-Pile is sourced from diverse, real-world long-video datasets and annotated through a carefully designed pipeline which ensures both quality and diversity.
-TimeScope is trained on ToTG-Pile using a two-stage supervised fine-tuning strategy.
-In the first stage, we warm up the model’s temporal grounding capability by directly predicting target time intervals from the input video and task description.
-In the second stage, we establish the progressive reasoning capability by training the model to first estimate coarse-grained temporal scopes using abstracted video representations, and then refine its predictions to identify fine-grained time intervals based on detailed representations.
-We conduct comprehensive experiments across a wide range of scenarios, evaluating TimeScope not only on ToTG-Bench but also on popular benchmarks for traditional temporal groundingCaba Heilbron et al.
-(2015); Sigurdsson et al.
-(2016); Cheng et al.
-(2025).
-The results demonstrate that TimeScope delivers significant improvements over existing approaches, including both dual-encoder-based model, specialized MLLMs for temporal grounding and advanced generic MLLMsBai et al.
-(2025); Yang et al.
-(2025a).
-Further analyses highlight the contributions of each component within our framework, validating its effectiveness in addressing this new and challenging problem.
-To facilitate future research, we will publicly release all resources, including model, dataset, benchmark, and source code.
+Multimodal large language models (MLLMs) have become increasingly prominent in tackling long-video understanding (LVU) problems.
+However, they still struggle with complex tasks which call for fine-grained details, especially those sparsely distributed across long videos.
+One promising strategy to mitigate this problem is to present MLLMs only with crucial temporal intervals which contains relevant information to their LVU tasks.
+However, existing temporal grounding (TG) methods are primarily designed for tasks with explicit descriptions of time-intervals, such as “the moments when the man is tripped by a stone and falls to the ground”, instead of directly handling native task requirements, e.g., “explain why the man in the video is sent to the hospital”.
+This gap prevents the direct utilization of existing TG methods for LVU tasks, resulting in a severe limitation for real-world applications.
+To formalize the above challenge, we define a new problem: Task-oriented Temporal Grounding (ToTG), where a model needs to localize crucial temporal intervals that are relevant to a specific downstream task based on its native requirement.
+For example, as illustrated in Figure 3, given the query “tell me what happens after cutting the vegetables,” the model identifies the interval corresponding to “stir-frying the vegetables.” Unlike traditional TG problems, ToTG is more closely aligned with practical applications, as it can be directly conducted to support real-world LVU tasks.
+The new ToTG problem introduces unprecedented technical challenges for existing methods on TG, as it requires both in-depth task comprehension and fine-grained temporal localization over long videos.
+In this paper, we present a systematic study of the ToTG problem.
+First, to address the absence of suitable resources for evaluating ToTG performance, we introduce a new benchmark, ToTG-Bench.
+ToTG-Bench incorporates 32 diverse video domains and 12 task categories, with video durations ranging from a few seconds to over an hour.
+Each instance is constructed through a human–machine collaborative annotation pipeline that ensures high-quality temporal localization.
+Together, these designs enable ToTG-Bench to provide a comprehensive evaluation of ToTG performance across a wide spectrum of task and video types.
+To further enhance the grounding model’s capability for ToTG, we develop a novel progressive-reasoning framework called TimeScope.
+As discussed, ToTG poses two major technical challenges: 1) in-depth understanding of the task, and 2) fine-grained temporal grounding within long videos.
+TimeScope addresses these challenges through two consecutive operations.
+First, the model focuses on comprehending the task and identifying which parts of the video are likely to contain the required information.
+To accomplish this, TimeScope performs chain-of-thought (CoT) reasoning based on the task description and a holistic abstraction of the video, producing a set of candidate time scopes that are likely to be relevant.
+Next, TimeScope further encodes the candidate scopes with more detailed visual information, and predicts the precise temporal intervals contained within them.
+This progressive reasoning pipeline not only improves the grounding precision but also enables more efficient processing of long videos.
+To obtain a general temporal-grounding capability across diverse long-video tasks, TimeScope is trained through extensive supervised fine-tuning.
+To construct the training data, we prompt a teacher model to generate chain-of-thought (CoT) reasoning for a wide range of LVU tasks.
+We then filter the generated CoT trajectories, retaining only those whose reasoning leads to precise temporal grounding.
+This process yields ToTG-Pile, a diverse and high-quality dataset tailored for ToTG.
+In our experiments, we evaluate TimeScope against traditional TG methods [1, 41] using not only the proposed ToTG-Bench, but also standard temporal-grounding benchmarks [2, 33, 6] and long-video understanding benchmarks [39, 47].
+Our evaluation demonstrates the effectiveness of TimeScope from three perspectives.
+First, it substantially improves temporal-grounding precision over existing TG baselines.
+Second, it exhibits strong generalizability, achieving consistent performance gains across heterogeneous benchmarks.
+Third, it significantly enhances the performance of downstream LVU tasks when used as a pre-localization module.
+Extended ablation studies further highlight the contribution of each component, reflecting the validity of our technical design.
+All resources, including the model, benchmark, dataset, and source code, will be publicly released to facilitate future research.
+The contributions of this paper are summarized as follows.
+1) We formulate the ToTG problem, which formally defines the localization of task-relevant information in long videos.
+2) We introduce ToTG-Bench, which enables unified and comprehensive evaluation of ToTG performance across diverse tasks and video domains.
+3) We propose TimeScope, a progressive reasoning framework for ToTG, and construct ToTG-Pile, a high-quality supervised fine-tuning dataset that enhances TimeScope’s capability.
+4) We conduct extensive experiments, demonstrating TimeScope’s empirical effectiveness in ToTG precision, generalization across benchmarks, and benefits to downstream LVU tasks.
+| Benchmark | Video Num. | Avg. Duration | Duration Range | Video Domain | Query Type |
+| Traditional Temporal Grounding Benchmark | |||||
+| Charades-STA | 1,331 | 29.9 s | 7.2 s - 1.2 min | Daily Activities | Explicit Description |
+| ActivityNet | 4,885 | 122.0 s | 2.3 s - 12.4 s | Daily Activities | Explicit Description |
+| V-STaR | 732 | 1.8 min | 15.0 s - 59.2 min | 9 Domains | Explicit Description |
+| Clue-grounded QA Benchmark | |||||
+| CG-Bench | 1,219 | 28.7 min | 9.1 min - 1.8 hr | 14 Domains | Perception Task |
+| Next-GQA (test) | 990 | 38.7 s | 10.0 s - 2.5 min | Daily Activities | Perception Task |
+| Task-oriented Temporal Grounding Benchmark | |||||
+| ToTG-Bench | 337 | 13.5 min | 30 s - 1.2 hr | 35 Domains | Perception/Reasoning Task |
 2 Related work
-2.1 Video Temporal Grounding
-The traditional temporal grounding (TG) task requires models to localize a time interval in a video given a query that explicitly describes the target event.
-Early approaches are mainly dual-encoder-based, where video and language features are extracted using different pre-trained encoders (e.g., BERT Devlin et al.
-(2019), CLIP Radford et al.
-(2021), SigLip Zhai et al.
-(2023)), and then fused for time interval prediction Mu et al.
-(2024); Lei et al.
-(2021); Moon et al.
-(2023b; a); Gordeev et al.
-(2024).
-These models rely on simple visual-text alignment, which limits their generalization to out-of-domain or more complex queries.
-More recently, researchers have explored applying MLLMs to temporal grounding Huang et al.
-(2024b); Ren et al.
-(2024); Zeng et al.
-(2025); Guo et al.
-(2024); Wang et al.
-(2025b); Li et al.
-(2025f).
-For instance, TimeChat Ren et al.
-(2024) introduces a time-aware frame encoder that binds visual tokens with their corresponding timestamps at the frame level for temporal grounding.
-Similarly, TimeSuite Zeng et al.
-(2025) proposes temporal-adaptive position encoding to strengthen temporal awareness in video representations.
-Trace Guo et al.
-(2024) designs a specialized encoder and head for timestamp input, while Time-R1 Wang et al.
-(2025b) employs a reasoning-guided post-training framework with reinforcement learning and verifiable rewards to improve grounding accuracy.
-In addition to these specialized MLLMs, recent generic MLLMs (e.g., Qwen2.5-VL Bai et al.
-(2025), Keye-VL-1.5 Yang et al.
-(2025b)) have also demonstrated certain capabilities for temporal grounding.
-Despite these advances, existing methods are still limited in handling complex queries that require not just locating explicit events but identifying intervals that support completing a task in long video.
-In particular, most MLLM-based approaches underutilize the generalization potential of MLLMs.
-Motivated by these limitations, we introduce the new problem of task-oriented temporal grounding, along with a benchmark, a dataset, and a dedicated framework to address it.
 2.1 Long Video Understanding
-The field of long video understanding (LVU) has developed rapidly in recent years, with many powerful MLLMs emerging, such as VideoChatFlash Li et al.
-(2025b), Video-XL-2 Qin et al.
-(2025a), Eagle2.5 Chen et al.
-(2025), and InternVL3 Zhu et al.
-(2025).
-These models demonstrate strong general video understanding capabilities and serve as versatile backbones for various video tasks, including temporal grounding.
-However, precisely localizing or perceiving fine-grained details within second-level intervals remains a major challenge for current LVU MLLMs.
-To address this, some works introduce additional modules to identify key frames or video segments based on task-oriented queries Wang et al.
-(2025a); Huang et al.
-(2025); Yu et al.
-(2024); Qin et al.
-(2025b).
+The field of long video understanding (LVU) has developed rapidly in recent years, with many powerful MLLMs emerging, such as VideoChatFlash [20], Video-XL-2 [29], Eagle2.5 [4], and InternVL3 [49].
+These models demonstrate strong general video understanding capabilities and serve as versatile backbones for various video tasks.
+However, precisely capturing fine-grained details within second-level intervals remains a major challenge for current LVU models.
+To address this, some works introduce additional modules to assist LVU models by identifying key frames [36, 15, 43, 30].
 These modules are typically similarity-based and thus lack deeper semantic understanding of the video content, limiting their compatibility with diverse downstream tasks in long video scenarios.
@@
 This enables the model to efficiently and accurately localize critical time intervals in long videos for a wide range of tasks.
+2.2 Video Temporal Grounding
+The traditional temporal grounding (TG) task requires models to localize a time interval in a video given a query that explicitly describes the target content.
+Early approaches are mainly dual-encoder-based, where video and language features are extracted using different pre-trained encoders (e.g., BERT [7], CLIP [31], SigLip [45]), and then fused for time interval xprediction [27, 17, 26, 25, 11, 35, 34].
+These models lack generalizability and can only be evaluated under few-shot setting across different benchmarks.
+More recently, researchers have explored using MLLMs for more general temporal grounding [14, 32, 44, 12, 38].
+For instance, TimeChat [32] introduces a time-aware frame encoder that binds visual tokens with their corresponding timestamps at the frame level for temporal grounding.
+Similarly, TimeSuite [44] proposes temporal-adaptive position encoding to strengthen temporal awareness in video representations.
+Trace [12] designs a specialized encoder and head for timestamp input, while Time-R1 [38] employs a reasoning-guided post-training framework with reinforcement learning and verifiable rewards to improve grounding accuracy.
+In addition to these specialized MLLMs, recent generic MLLMs (e.g., Qwen2.5-VL [1], Keye-VL-1.5 [42]) have also demonstrated certain capabilities for temporal grounding.
+However, localizing small intervals in long videos is challenging due to their limited context window.
+UniTime [22] tackles this by adjusting the frame sampling rate for multi-stage grounding, but its sparse frame sampling can disrupt continuous event semantics, compromising precise temporal grounding.
+In addition, previous approaches tend to fall short when it comes to more complex and practical grounding tasks.
+Motivated by these limitations, we introduce the new problem of task-oriented temporal grounding, along with a benchmark, a dataset, and a dedicated framework to address it.
+3 Problem Definition
+We formally define the Task-oriented Temporal Grounding (ToTG) problem in this section.
+Formally, the input consists of a long video and a task-oriented natural language query .
+ToTG assumes that the critical visual information required to answer the task is not explicitly described in .
+Therefore, the model must localize the temporal interval containing this task-critical but implicit evidence without relying on direct semantic matches to the query.
+Let the target interval be denoted as .
+We define ToTG as:
+where is a grounding function that selects the video segment containing the necessary task-related visual cues, even though such cues are not directly mentioned in the query.
+This formulation represents a broad variety of real-world scenarios (such as instructional videos, plot reasoning, and surgical workflows) where the required visual evidence is implicit and must be located in the video rather than retrieved through explicit semantic matching.
 4 Method
-In this section, we introduce our new proposed problem, Task-Oriented Temporal Grounding (ToTG) along with its associated benchmark ToTG-Bench, grounding framework TimeScope and training dataset ToTG-Pile in detail.
-In the following, Section 3.1 formulates the ToTG, which is a more challenging and valuable task compared to traditional temporal grounding.
-Section 3.2 presents ToTG-Bench, a comprehensive benchmark designed to evaluate model in localizing time intervals based on task description across various long video scenarios.
-Section 3.3 details our novel framework TimeScope, which aims to address the challenges of TOTG via progressive reasoning.
-Section 3.4 introduces the dataset TOTG-Piles and its carefully designed pipeline.
-3.1 Problem Formulation
-We formally define the Task-oriented Temporal Grounding (ToTG) problem in this section.
-Given an untrimmed video and a task-oriented query , the goal is to localize a temporal interval within that contains the information necessary to accomplish the task described by .
-In conventional temporal grounding, the query is usually an event description, including explicit information of the grounding target.
-This enables the model to locate the corresponding time interval via relatively straightforward visual-text alignment.
-In contrast, the query in ToTG supplements a higher-level task instruction about the target on this basis.
-Therefore, the model must locate the significant time interval by comprehensively understanding both visual and textual content, rather than relying solely on surface-level alignment.
-3.2 Benchmarks
-To facilitate the evaluation of ToTG, we introduce ToTG-Bench, a comprehensive and challenging benchmark.
-Unlike conventional temporal grounding benchmarks Sigurdsson et al.
-(2016); Caba Heilbron et al.
-(2015), ToTG-Bench features queries spanning 12 distinct task types.
-Its videos are drawn from diverse long-video understanding scenarios and cover durations ranging from a few seconds to nearly one hour.
-We construct the benchmark through a combination of an effective data filtering pipeline and careful manual annotation, ensuring accurate temporal intervals for each test sample.
-All of these make ToTG-Bench a high-quality, diverse, and challenging testbed for advancing research in ToTG.
-Benchmark Construction We first collect samples from four long-video understanding benchmarks (MLVU, Video-MME, LongVideoBench, and V-STaR) as the candidate data for ToTG-Bench construction.
-The questions in these samples are used as task-oriented queries for grounding.
-However, not all data are suitable for ToTG, so we design the following filtering pipeline to obtain qualified candidates: 1) Task type filtering.
-We exclude samples whose task types are not compatible with ToTG (e.g., multilingual, summarization, event-ordering tasks and etc.).
-2) Uniqueness filtering.
-We retain only samples where the query corresponds to a unique single grounding target.
-Specifically, we divide each long video into segments and evaluate them using a temporal grounding model followed by manual verification.
-Samples with exactly only one valid interval contained target are retained.
-3) Information validation.
-We further verify that the preserved intervals contain sufficient information to answer the query.
-We use a advanced MLLM to generate answers based on sampled frames from the predicted intervals, and discard samples for which the answers fail.
-This pipeline yields a set of qualified candidate data with unique grounding targets, ensuring both annotation quality and fairness in subsequent evaluation.
-Furthermore, we balance the distribution of candidate data across query task types, video categories, video durations, and temporal positions of the target intervals (We refer to this as Query Center) so that each time interval within videos has a roughly equal chance of containing the target.
-Finally, we carefully manual annotate time interval to ensure the quality of each sample.
-Benchmark Statistics As illustrated in Figure 2, ToTG-Bench demonstrates significant diversity across task types, video categories, video durations and position of target interval.
-Specifically, it includes 12 task types (e.g., action reasoning, OCR perception temporal reasoning and etc.), covering a broad range of video understanding tasks 111We follow the definition of task type and video category from Video-MME.
-It further comprises 35 video categories, covering a wide range of real-world video domains.
-The video durations in ToTG-Bench range from a few minutes to up to one hour.
-Moreover, the target intervals are balanced to be uniformly distributed along the video timeline, ensuring unbiased evaluation.
+This section outlines the benchmark ToTG-Bench, the proposed framework TimeScope, and the specialized dataset ToTG-Pile in Sec. 4.1, Sec. 4.2, and Sec. 4.3, respectively.
+4.1 ToTG-Bench
+To enable a comprehensive evaluation of ToTG, we introduce ToTG-Bench.
+We construct ToTG-Bench based on long-video QA datasets, since their most questions naturally align with the definition of ToTG—queries that implicitly specify task-required information within extended video contexts.
+Moreover, such datasets cover a wide range of domains and task types, allowing us to build a benchmark that is both diverse and realistic.
+Concretely, we collect question–answer pairs from four public long-video understanding datasets [47, 10, 39, 5], using the questions as task instruction style queries and deriving the corresponding ground-truth temporal intervals through a semi-automated annotation pipeline.
+First, Gemini-2.5-Pro is used to predict multiple temporal intervals as candidate segments.
+These candidates are then carefully verified, selected, and refined manually to produce the final grounded annotations.
+As shown in Table 1, ToTG-Bench exhibits substantially higher diversity and comprehensiveness compared with both traditional temporal grounding benchmarks (Charades-STA [33], ActivityNet [2], and V-STaR [5]) and clue-grounded QA benchmarks (CG-Bench [3] and Next-GQA [40]).
+It covers a wide range of video durations—from seconds to over one hour—with an average length of 13.5 minutes, and spans 35 realistic video domains (e.g.
+vlogs, news, documentaries, and sports).
+Moreover, ToTG-Bench incorporates from both perception tasks (e.g., action or object recognition) and reasoning tasks (e.g., temporal or causal reasoning), covering a total of 12 distinct task types.
+These characteristics make ToTG-Bench a diverse, comprehensive, and realistic foundation for benchmarking task-oriented temporal grounding models.
+More details about ToTG-Bench can be found Appendix.
 4.2 TimeScope
-ToTG requires models to identify the necessary information needed to accomplish diverse tasks in long videos.
-However, existing temporal grounding models are often distracted by abundant irrelevant content in long contexts, which prevents them from accurately grounding complex task-oriented queries.
-To overcome this, we design TimeScope, a progressive reasoning framework that adapts MLLM backbones for task-oriented temporal grounding.
-Instead of directly predicting over the full long video at once, TimeScope performs reasoning in two complementary stages, progressively refining from holistic context to precise details (Figure 1).
-Stage 1 (Coarse reasoning).
-TimeScope first compresses the cached Key-Value (KV) states of visual tokens (detonated as ) obtained during pre-filling.
-Through average pooling, the fine-grained are distilled into , which serves as a compact representation of the global video context.
-On this compact KV space, the model performs a holistic reasoning step to hypothesize a coarse temporal window that is likely to contain the target.
-Stage 2 (Fine reasoning).
-Given the temporal window , TimeScope selectively reloads only the fine-grained from cache for the corresponding frames, while discarding irrelevant context outside .
-This selective refinement enables the model to reason over detailed visual cues within the localized window, and to output a precise grounding interval.
-By framing temporal grounding on long frames as a progressive reasoning process, TimeScope achieves both superior efficiency and accuracy in task-oriented long video temporal grounding.
-3.4 Datasets
-Existing temporal grounding datasets are mostly limited to short videos with explicit event queries, which makes models trained on them poorly suited for ToTG task.
-To bridge this gap, we curate a new large-scale dataset, ToTG-Pile, specifically designed for task-oriented temporal grounding.
-ToTG-Pile is diverse and comprehensive, incorporating both traditional temporal grounding data and newly constructed task-oriented data.
-Traditional temporal grounding data.
-Existing datasets for temporal grounding are limited in terms of video duration.
-To address this, we recaption existing short-video datasets using Qwen2.5-VL to generate concise descriptions.
-We then concatenate these short videos into longer videos, and the concise descriptions are used to construct explicit temporal grounding queries.
-Through this pipeline, we obtain a training set of 85k long-video temporal grounding samples, with an average video duration of about 500 seconds (approximately 8 minutes).
-Task-oriented temporal grounding data.
-For ToTG data, we collect raw samples from NextQA Xiao et al.
-(2021), STAR Wu et al.
-(2024), and CLEVRER Yi et al.
-(2020).
-As illustrated in Figure 1(b), the video and answer of each sample are fed into multiple temporal grounding models (refer to Video Expert) to generate candidate time intervals containing the answer.
-A sample is retained only if the Intersection-over-Union (IoU) between intervals predicted by different models exceeds 0.5, and is further evaluated by another MLLM to ensure it contains sufficient information to answer the question.
-The resulting qualified QA pairs with timestamps are then concatenated into longer videos, yielding training data tailored for task-oriented grounding.
-Overall, ToTG-Pile unifies traditional grounding data with newly constructed task-oriented data, ensuring diversity across tasks, durations, and video domains.
-ToTG-Pile lays the foundation for developing excellent temporal grounding models in ToTG task.
-4 Experiment
-4.1 Implementation details
-We adopt VideoXL-2 as our backbone for two reasons: first, it can ingest very long video sequences, which makes it straightforward to build long-video temporal-understanding methods; second, its internal design interleaves timestamp tokens, giving the model a strong built-in sense of time.
-TimeScope is trained on ToTG-Pile with a two-stage supervised fine-tuning schedule.
-In Stage 1 we use the temporal-grounding splits of ToTG-Pile and ask the model to predict the target time interval directly from the raw video and the task description, thereby bootstrapping its basic localization ability.
-In Stage 2 we apply heavy temporal augmentations—random cropping, shifting, and scaling of the time span—to the training videos, forcing the model to first estimate a coarse temporal window from an abstract video representation and then refine that window into a fine-grained interval using the detailed representation.
-Throughout training we sample video frames at 1 fps, capping at 300 frames maximum.
+We propose TimeScope, a progressive reasoning framework that reduces the search space in a coarse-to-fine manner: it first identifies a coarse temporal interval where the target is likely to occur, and then performs refined grounding within this narrowed region.
+Holistic & Detailed Representation.
+Dense frame sampling is crucial for accurate grounding in long videos, as it preserves fine-grained temporal continuity and reduces the risk of missing brief task-critical moments.
+However, existing methods such as UniTime [22] are limited to sparse sampling due to the prohibitive memory cost of processing long sequences.
+To reconcile dense sampling with memory constraints, TimeScope decouples global context from local details.
+Specifically, we define a Holistic Video Representation (HVR) to abstract long-range context with minimal memory overhead, and a Detailed Video Representation (DVR) to retain high-resolution visual cues for precise localization.
+To avoid the memory overflow of processing all frames simultaneously, TimeScope constructs both representations in a streaming manner.
+As illustrated in Fig.
+2 (right), the densely sampled video is divided into temporal segments .
+Each segment is sequentially processed by the MLLM to produce both representations.
+For each layer , the key–value states computed from form its detailed representation .
+These KVs are then temporally downsampled to obtain the holistic counterpart , which summarizes segment-level semantics at significantly reduced resolution.
+After that, all previously generated holistic caches are concatenated to form a lightweight historical memory.
+The current detailed representation attends to this holistic memory through cross-attention, enabling efficient integration of long-range temporal information without incurring full computation on the entire video.
+The resulting updated states are forwarded to obtain .
+After all segments are processed, the per-segment representations and are concatenated along the temporal dimension to form the final detailed- and holistic-level summaries of the entire video, which are stored for subsequent progressive reasoning.
+| Method | S(R1@0.3) | S(R1@0.5) | M(R1@0.3) | M(R1@0.5) | L(R1@0.3) | L(R1@0.5) |
+| Temporal Grounding Models | ||||||
+| TimeR1-7B [38] | 20.1 | 16.0 | 6.8 | 4.1 | 15.8 | 11.0 |
+| VideoChatR1-7B [21] | 17.3 | 14.0 | 6.6 | 5.6 | 12.1 | 10.9 |
+| Temporal-RLT-7B [18] | 54.0 | 35.3 | 17.9 | 12.3 | 14.1 | 11.5 |
+| UniTime [23] | 52.4 | 42.7 | 29.6 | 23.2 | 24.2 | 22.6 |
+| Video Understanding Models | ||||||
+| Qwen-2.5VL-7B [1] | 48.2 | 40.6 | 12.2 | 10.3 | 14.6 | 12.3 |
+| Keye-VL-1.5-8B [41] | 36.0 | 34.7 | 25.4 | 25.4 | 32.7 | 28.6 |
+| TimeScope-7B | 52.3 | 46.3 | 45.2 | 42.3 | 47.3 | 37.8 |
+| Model | OR | SR | TR | AR | SP | AP | TP | EG | OP | OB | AT | CP | TU |
+| TimeR1-7B | 13.8 | 17.2 | 11.8 | 6.50 | 2.90 | 14.8 | 8.80 | 8.20 | 7.70 | 17.4 | 7.70 | 13.3 | 12.5 |
+| Temporal-RLT-7B | 22.4 | 20.7 | 23.5 | 8.70 | 42.9 | 48.1 | 23.5 | 6.40 | 26.9 | 23.9 | 10.8 | 40.0 | 6.20 |
+| Qwen-2.5VL-7B | 24.1 | 34.5 | 23.5 | 19.6 | 1.20 | 40.7 | 17.6 | 10.0 | 30.8 | 28.3 | 18.7 | 33.3 | 9.30 |
+| Keye-VL-1.5-8B | 19.0 | 27.6 | 35.3 | 41.3 | 10.7 | 22.2 | 55.9 | 10.0 | 26.9 | 30.4 | 36.9 | 26.7 | 56.2 |
+| TimeScope-7B | 42.9 | 48.3 | 47.1 | 47.8 | 36.7 | 63.0 | 35.3 | 30.0 | 46.2 | 42.9 | 36.9 | 40.0 | 35.8 |
 | Method | Charades-STA | ActivityNet | ||||
@@
 | Open-source VLP Method | ||||||
-| 2D-TAN Zhang et al. (2020) | 45.8 | 27.9 | – | 60.4 | 43.4 | – |
-| UniVTG Lin et al. (2023) | 60.2 | 38.6 | – | 56.1 | 43.4 | – |
-| SSRN Zhu et al. (2022) | 65.5 | 42.6 | – | – | 54.5 | – |
-| SnAG Mu et al. (2024) | 64.6 | 46.2 | – | – | 48.6 | – |
-| EaTR Jang et al. (2023) | 68.4 | 44.9 | – | – | 58.2 | – |
+| 2D-TAN [46] | 45.8 | 27.9 | – | 60.4 | 43.4 | – |
+| UniVTG [24] | 60.2 | 38.6 | – | 56.1 | 43.4 | – |
+| SSRN [48] | 65.5 | 42.6 | – | – | 54.5 | – |
+| SnAG [27] | 64.6 | 46.2 | – | – | 48.6 | – |
+| EaTR [16] | 68.4 | 44.9 | – | – | 58.2 | – |
 | Open-source MLLMs Method | ||||||
-| TimeChat Ren et al. (2024) | 32.2 | 13.4 | 32.2 | 36.2 | 20.2 | 21.8 |
-| VTimeLLM Huang et al. (2024a) | 27.5 | 11.4 | 31.2 | 44.0 | 27.8 | 30.4 |
-| VideoChat-Flash Li et al. (2024) | 53.1 | 27.6 | – | – | – | – |
-| TRACE Guo et al. (2024) | 61.7 | 41.4 | 41.4 | 37.7 | 24.0 | 39.0 |
-| HawkEye Wang et al. (2024) | 58.3 | 28.8 | – | 55.9 | 34.7 | – |
-| TimeSuite Zeng et al. (2025) | 67.1 | 43.0 | – | – | – | – |
-| Time-R1 Wang et al. (2025b) | 72.2 | 50.1 | – | 58.6 | 39.0 | – |
-| DeepVideo-R1-7B Park et al. (2025) | 71.7 | 50.6 | 61.2 | 33.9 | 18.0 | 36.9 |
-| VideoChat-R1-7B Li et al. (2025c) | 71.7 | 50.2 | 60.8 | 33.4 | 17.7 | 36.6 |
-| TimeZero-7B Wang et al. (2025b) | 60.8 | 35.3 | 58.1 | 39.0 | 21.4 | 40.5 |
-| Temporal-RLT-7BLi et al. (2025a) | 67.9 | 44.1 | 57.0 | 38.4 | 20.2 | 39.0 |
-| TimeScope-7B | 78.9 | 64.0 | 57.5 | 69.6 | 59.0 | 48.0 |
-4.2 Benchmarks and Metrics
-For evaluation, we benchmark our method across the following four categories: (i) short-video temporal grounding, including Charades-STA [46] and ActivityNet-Captions; (ii) long-video temporal grounding, including videos longer than 300s from V-STaR and Vid-Chapters-7M; (iii) Video Question Answering (VideoQA), including three general long-video QA benchmarks: CG-Bench, MLVU, and LongVideoBench ; (iv) long video task grounding, including our proposed ToTG-bench.
-The statistics of the evaluation benchmarks used are listed in Table 5.
-Evaluation metrics.
-For temporal grounding tasks and task grounding tasks, we adopt Recall@1 (R@1) at multiple temporal intersection-over-union (IoU) thresholds and mean IoU (mIoU) as evaluation metrics.
-Specifically, for temporal grounding benchmarks, we use IoU thresholds of 0.3, 0.5, and 0.7.
-For general VideoQA tasks, we report standard accuracy metrics.
-| Model | R1@0.5 | R1@0.7 |
-| VTimeLLM-7B Huang et al. (2024a) | 4.1 | 1.6 |
-| CLIP Radford et al. (2021) | 5.2 | 2.3 |
-| M-DETR Kamath et al. (2021) | 27.3 | 17.6 |
-| ReVisionLLM Hannan et al. (2024) | 27.4 | 21.8 |
-| Qwen2.5-VL-7B Bai et al. (2025) | 0.3 | 0.1 |
-| VITAL-7B Zhang et al. (2025) | 25.8 | 19.5 |
-| TimeScope-7B | 24.3 | 22.1 |
+| TimeChat [32] | 32.2 | 13.4 | 32.2 | 36.2 | 20.2 | 21.8 |
+| VTimeLLM [13] | 27.5 | 11.4 | 31.2 | 44.0 | 27.8 | 30.4 |
+| VideoChat-Flash [19] | 53.1 | 27.6 | – | – | – | – |
+| TRACE [12] | 61.7 | 41.4 | 41.4 | 37.7 | 24.0 | 39.0 |
+| HawkEye [37] | 58.3 | 28.8 | – | 55.9 | 34.7 | – |
+| TimeSuite [44] | 67.1 | 43.0 | – | – | – | – |
+| Time-R1 [38] | 72.2 | 50.1 | – | 58.6 | 39.0 | – |
+| DeepVideo-R1-7B [28] | 71.7 | 50.6 | 61.2 | 33.9 | 18.0 | 36.9 |
+| VideoChat-R1-7B [21] | 71.7 | 50.2 | 60.8 | 33.4 | 17.7 | 36.6 |
+| TimeZero-7B [38] | 60.8 | 35.3 | 58.1 | 39.0 | 21.4 | 40.5 |
+| Temporal-RLT-7B[18] | 67.9 | 44.1 | 57.0 | 38.4 | 20.2 | 39.0 |
+| TimeScope-7B | 78.9 | 61.2 | 56.2 | 66.9 | 56.0 | 46.0 |
+Progressive Reasoning.
+With both and prepared, TimeScope performs grounding in a coarse-to-fine manner.
+In the first step, only the compact holistic representation is kept in GPU memory, while the high-resolution remains stored in CPU memory.
+Leveraging the lightweight long-range context encoded in , the MLLM efficiently predicts a coarse temporal interval that is most likely to contain the task-relevant moment for the query .
+This stage effectively eliminates the majority of irrelevant frames and substantially reduces the temporal search space.
+Next, TimeScope reloads only the fine-grained corresponding to the predicted interval and performs detailed reasoning.
+The rich local temporal and visual details preserved in allow the model to refine the boundaries within the coarse region and accurately localize the target moment.
+By combining the global efficiency of with the local precision of , the progressive reasoning process achieves high localization accuracy while maintaining low computational cost of processing the long video at high resolution.
+4.3 ToTG-Pile
+To maximize TimeScope’s capacity for task-oriented temporal grounding, we created ToTG-Pile—a large-scale dataset bridging conventional temporal grounding data with Task-oriented Temporal Grounding requirements.
+ToTG-Pile features two key characteristics: (1) all samples follow task-oriented query definitions (Sec.
+3), where queries are formulated as task instructions rather than explicit temporal descriptions, and (2) each sample includes both ground-truth temporal intervals and chain-of-thought (CoT) annotations that detail the localization reasoning process.
+We constructed the dataset by collecting diverse videos with broad coverage of tasks and visual contexts.
+Specifically, we leverage comprehensive Video VQA datasets such as VideoR1 [9] and FineVideo [8].
+To curate high-quality reasoning data, we employ a three-stage pipeline: (1) Answer-aware prompting: we feed answers from original VQA samples into expert temporal grounding models to obtain candidate temporal intervals; (2) Cross-validation filtering: we leverage multiple expert models for cross-validation, discarding low-quality samples with IoU 0.1 across models; (3) CoT annotation generation: we use reasoning-capable MLLMs to generate chain-of-thought annotations that detail the localization reasoning process.
+Additionally, to extend the dataset to long-video scenarios, we synthesized 90K long clips (10 minutes each) by concatenating short videos and generated corresponding temporal grounding queries from video captions.
+This augmentation enhances TimeScope’s capability in handling both conventional temporal grounding and task-oriented temporal grounding in extended video contexts.
+Overall, ToTG-Pile provides a large-scale, diverse, and reasoning-oriented foundation for training MLLMs to perform effective and generalizable task-oriented temporal grounding in long videos.
+5 Experiment
+5.1 Implementation details
+We adopt VideoXL-2 [29] as our backbone for two reasons: (1) it can process very long video sequences, enabling straightforward construction of long-video temporal understanding methods, and (2) its internal design interleaves timestamp tokens, providing the model with a strong built-in temporal awareness.
+Stage 1: Basic Localization.
+We use the temporal grounding splits of ToTG-Pile and train the model to predict target time intervals directly from raw video and task descriptions, bootstrapping its basic localization ability.
+Stage 2: Coarse-to-Fine Refinement.
+We apply heavy temporal augmentations (random cropping, shifting, and scaling of time spans) to training videos, forcing the model to first estimate a coarse temporal window from abstract video representations and then refine it into fine-grained intervals using detailed representations.
+5.2 Results on Task-oriented temporal grounding
+Evaluation on ToTG-Bench.
+We evaluate TimeScope on ToTG-Bench, which categorizes videos into three duration segments: short (180s), medium (180–600s), and long (600s).
+We report IoU@0.3 and IoU@0.5 metrics and compare TimeScope against both specialized temporal understanding MLLMs and general video understanding models.
+As shown in Table 2, TimeScope demonstrates strong and consistent performance across all video durations.
+Most notably, TimeScope outperforms all baselines by 20–30 points on medium and long videos, demonstrating substantial advantages in handling task-oriented temporal grounding in extended contexts.
+On short videos, TimeScope achieves competitive performance, trailing the state-of-the-art Temporal-RLT [18] by less than 2 points while significantly surpassing it on longer videos.
+These results highlight TimeScope’s effectiveness in task-oriented temporal localization, particularly its ability to perform accurate reasoning and grounding in long-video scenarios while maintaining robust performance across different video durations.
+We further evaluate TimeScope and baseline models on ToTG-Bench across different task categories using IoU@0.5 metric, as shown in Table 3.
+TimeScope demonstrates strong performance across all task categories.
+Notably, in reasoning tasks (including Object Reasoning, Spatial Reasoning, Temporal Reasoning, and Action Reasoning), TimeScope outperforms all baselines by 10–20 points, highlighting its superior capability in task-oriented reasoning and temporal localization.
+Across other task categories, TimeScope consistently achieves top-tier performance, demonstrating robust generalization to diverse task types.
 | Model | R1@0.5 | R1@0.7 |
 | Qwen2.5-VL-7B | 0.0 | 0.0 |
-| UniTime | 0.629 | 0.629 |
-| Keye-1.5-VL-8B | 0.639 | 0.491 |
-| TimeScope-7B | 0.909 | 0.909 |
+| UniTime | 62.9 | 62.9 |
+| Keye-1.5-VL-8B | 63.9 | 49.1 |
+| TimeScope-7B | 87.5 | 85.2 |
+5.3 Results on Traditional Temporal Grounding
+We conduct a comprehensive comparison of TimeScope against traditional and MLLM-based methods on conventional temporal grounding benchmarks, covering both short-video and long-video settings.
+Short-video Benchmarks.
+As shown in Table 4, TimeScope achieves state-of-the-art performance across all short-video benchmarks.
+On Charades-STA, TimeScope attains an R1@0.7 score of 64.0, significantly surpassing VideoChat-Flash (27.6), TimeSuite (43.0), and Time-R1 (50.1).
+On ActivityNet, it achieves an R1@0.7 score of 59.0, outperforming HawkEye (34.7) and Time-R1 (39.0).
+Notably, TimeScope maintains a smaller gap between R1@0.5 and R1@0.7 compared to most baselines, indicating its capability for more precise temporal localization.
+Long-video Benchmarks.
+As shown in Table 5, TimeScope achieves an R1@0.7 score of 85.2 on V-STaR (with videos up to 300 seconds), substantially exceeding UniTime (62.9) and Keye-1.5-VL (49.1).
+This demonstrates TimeScope’s strong capability in handling extended temporal grounding tasks.
 | Method | CG-Bench | MLVU | LongVideoBench |
@@
 | Uniform Sample | 33.87 | 60.53 | 54.82 |
-| UniVTG Lin et al. (2023) | 34.87 | 62.56 | 54.67 |
-| VTimeLLM Huang et al. (2024a) | 34.60 | 59.52 | 54.30 |
-| TimeSuite Zeng et al. (2025) | 32.47 | 58.51 | 53.25 |
-| UniTime-Full Li et al. (2025f) | 40.30 | 66.50 | 56.47 |
-| TimeScope-7B | 39.53 | 70.14 | 59.18 |
-| Method | S(R1@0.5) | S(R1@0.7) | M(R1@0.5) | M(R1@0.7) | L(R1@0.5) | L(R1@0.7) |
-| no option | ||||||
-| Qwen-2.5VL-7B | 0.458 | 0.400 | 0.111 | 0.091 | 0.030 | 0.030 |
-| ARC-Hunyuan-Video-7B | 0.523 | 0.484 | 0.36 | 0.260 | 0.175 | 0.162 |
-| Keye-VL-1.5-8B | 0.521 | 0.333 | 0.313 | 0.261 | 0.221 | 0.124 |
-| UniTime | 0.524 | 0.427 | 0.296 | 0.232 | 0.242 | 0.226 |
-| TimeScope-7B | 0.574 | 0.436 | 0.47 | 0.44 | 0.435 | 0.405 |
-| has option | ||||||
-| Qwen-2.5VL-7B | 0.475 | 0.413 | 0.111 | 0.091 | 0.030 | 0.030 |
-| ARC-Hunyuan-Video-7B | 0.158 | 0.148 | 0.111 | 0.090 | 0.124 | 0.092 |
-| Keye-1.5-VL-8B | 0.527 | 0.363 | 0.514 | 0.471 | 0.261 | 0.185 |
-| UniTime | 0.557 | 0.441 | 0.317 | 0.274 | 0.287 | 0.256 |
-| TimeScope-7B | 0.683 | 0.515 | 0.52 | 0.44 | 0.467 | 0.426 |
-4.3 SOTA Performance on Traditional Temporal Grounding
-We conduct a comprehensive comparison of TimeScope against dual-encoder-based methods, MLLM-based methods, and video understanding models on both short- and long-video temporal grounding benchmarks.
-As shown in Table 1, TimeScope achieves state-of-the-art (SOTA) performance across all benchmarks.
-For instance, on Charades-STA, TimeScope reached an R1@0.7 score of 64.0, surpassing VideoChat-Flash (27.6), TimeSuite (43.0), and Time-R1 (50.1).
-On ActivityNet, it achieved an R1@0.7 score of 59, outperforming HawkEye (34.7) and Time-R1 (39.0).
-Particularly noteworthy is that, compared to most other models, TimeScope maintains a smaller gap between R@1@0.5 and R@1@0.7.
-This indicates that our model is capable of more precise temporal localization.
-Furthermore, as shown in Table 3, TimeScope achieved an R1@0.7 score of 90.9 on V-STaR, where video duration reaches 300 seconds, significantly surpassing UniTime’s 62.9 and Keye-1.5-VL’s 49.1, demonstrating its superior performance and strong capability in handling long-video grounding tasks.
-4.4 SOTA Performance on Task-oriented temporal grounding
-Besides traditional temporal grounding, we also evaluate TimeScope on ToTG-Bench, comparing it with MLLM-based models and generic MLLMs.
-As shown in Table LABEL:lvtgbench, where “has option” indicates that the options are fed into the model along with certain prompt information, TimeScope achieves outstanding and robust performance across different video durations.
-Specifically, while TimeScope and ARC-Hunyuan-Video-7B Ge et al.
-(2025) achieve comparable performance on short videos (within a 5-point difference), TimeScope shows clear superiority on medium and long videos, where it outperforms other models by 10 to 20 points.
-This indicates that our approach can effectively handle videos of various durations, especially long videos, and demonstrates excellent grounding ability for task-oriented queries.
-| Model | Early | Middle | Late | Std Dev () |
-| Qwen-2.5-VL Bai et al. (2025) | 20.0 | 13.4 | 4.1 | 8.0 |
-| Keye-1.5-VL Yang et al. (2025a) | 50.0 | 33.3 | 31.1 | 10.4 |
-| UniTime-Full Yang et al. (2025a) | 50.9 | 28.1 | 44.0 | 11.6 |
-| TimeScope-7B | 39.3 | 44.8 | 46.0 | 3.5 |
-4.5 Improvement for Long-Video QA
-As discussed in Section 3.1, the strong performance of TimeScope on the ToTG task suggests that it can help MLLMs capture the critical information in long videos required for question answering.
-To validate this point, we first use TimeScope and other temporal grounding models to localize time interval.
-The frames within the predicted intervals are then fed into Qwen2.5-VL-7B for answering.
-Their results are compared against the default setting, where frames are uniformly sampled without temporal grounding.
-We evaluate this experiment on three long-video understanding benchmarks: CG-Bench, MLVU, and LongVideoBench.
-As shown in Table 4(a), TimeScope consistently brings significant improvements across all benchmarks, while other temporal grounding models fail to outperform the default uniform-sampling baseline.
-| Model | R1@0.3 | R1@0.5 | R1@0.7 |
-| TimeScopeScope | |||
-| Zeroshot | 89.7 | 85.2 | 73.8 |
-| TimeScopeScope | |||
-| Multi-stage | 92.1 | 90.9 | 90.9 |
-| Method | Prefill | Decode | Cost |
-| zero shot | |||
-| Stage1 | 1400ms | 672ms | 2072ms |
-| Multi-stage | |||
-| Stage1 | 1400ms | 441ms | |
-| Stage2 | 158ms | 357ms | 2356ms |
+| UniVTG [24] | 34.87 | 62.56 | 54.67 |
+| VTimeLLM [13] | 34.60 | 59.52 | 54.30 |
+| TimeSuite [44] | 32.47 | 58.51 | 53.25 |
+| UniTime-Full [23] | 40.30 | 66.50 | 56.47 |
+| TimeScope-7B | 38.47 | 68.12 | 58.34 |
+5.4 Benefits to Long-Video Understanding
+As discussed in Section 3, TimeScope’s strong performance on task-oriented temporal grounding demonstrates its potential to help MLLMs capture critical information in long videos for question answering.
+To validate this, we conduct experiments where TimeScope and baseline temporal grounding models first localize relevant time intervals, and then feed frames from the predicted intervals into Qwen2-VL-7B [1] for answer generation.
+We compare these results against other grounding models and a default uniform sampling baseline without temporal grounding.
+We evaluate on three long-video understanding benchmarks: CG-Bench, MLVU, and LongVideoBench.
+As shown in Table 6, TimeScope demonstrates strong performance and brings substantial improvements over uniform sampling across all benchmarks, surpassing most temporal grounding baselines.
+On MLVU and LongVideoBench, TimeScope achieves the highest scores of 68.12 and 58.34 respectively, demonstrating its effectiveness in identifying task-relevant temporal segments for long-video understanding.
+Notably, while many video understanding questions require information beyond a single temporal segment, TimeScope’s task-oriented grounding still provides meaningful performance gains, validating its practical utility for downstream applications.
 5.5 Ablation Studies
 Effectiveness of Progressive Reasoning.
-To assess the effectiveness and necessity of progressive reasoning, we conduct an ablation study comparing two settings: progressive reasoning and single-step reasoning.
-The study is performed on videos longer than 300 seconds from the V-STaR benchmark, with results reported in Table 4(a).
-We observe that progressive reasoning achieves substantial improvements over the single-step baseline.
-In particular, for longer videos, progressively narrowing the search space proves both effective and necessary for accurate temporal grounding.
+To evaluate the effectiveness and necessity of progressive reasoning in long-video scenarios, we conducted an ablation study comparing two settings: progressive reasoning versus standard single prediction.
+The experiment was performed on samples longer than 300 seconds from the V-STaR benchmark and samples exceeding 600 seconds from the ToTG-Bench, with results presented in Table 7.
+The findings demonstrate that progressive reasoning achieves substantial improvements over the single-step baseline across both traditional temporal grounding tasks and task-oriented tasks in long videos.
+Particularly notable is the performance gain huge improvement in IoU@0.7, which confirms the effectiveness of progressively narrowing the search space for achieving precise temporal localization.
 Efficiency of Progressive Reasoning.
-In addition to effectiveness, we also evaluate the computational efficiency of TimeScope.
-As shown in Table 4(b), progressive reasoning introduces only marginal additional overhead compared to single-step reasoning, while still delivering substantial accuracy gains.
-This efficiency largely benefits from our design of adjusting KV sparsity during progressive reasoning.
-Overall, TimeScope strikes a cost-effective balance between accuracy and computational cost, making it practical for long-video temporal grounding.
-Robustness against Time Bias.
-After analyzing the evaluation results on ToTG-Bench, we find that the Query Center (the position of the target time interval) introduces a significant bias that strongly influences the performance of temporal grounding models.
-As shown in Table 6, Qwen2.5-VL and Keye-1.5-VL tend to perform better when the target time interval appears at the beginning of the video, but their performance drops substantially when the target lies in the middle or towards the end.
-In contrast, our TimeScope model maintains robust performance across different target interval positions, demonstrating its resilience against this bias.
+Despite employing progressive reasoning to achieve higher prediction accuracy, TimeScope remains highly efficient—thanks to its streaming-based video representation construction and carefully designed dual-level representation.
+As shown in Table 8, our framework enables the MLLM backbone to process significantly more frames while simultaneously improving throughput in both the prefill and decode stages, outperforming baseline methods without TimeScope in both speed and scalability.
+| Method | V-STaR(long) | ToTG-bench(long) | ||
+| R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
+| Zeroshot | 73.2 | 61.4 | 37.8 | 29.5 |
+| Progressive Reasoning | 86.4 | 83.0 | 37.8 | 34.0 |
+Effectiveness of Holistic Video Representation.
+As outlined in Sec.
+4.2, TimeScope enables dense frame sampling for long videos—made feasible by the efficient design of its Holistic Video Representation (HVR).
+As demonstrated in Table 9(a), with HVR, TimeScope supports dense sampling even for extremely long videos (up to 2,000 frames).
+This capability directly translates into significant gains in temporal grounding precision, particularly for fine-grained event localization.
+| Method | Frames | Prefill | Decode | Sum |
+| w/o TimeScope | 800 | 2903ms | 816ms | 3719ms |
+| w TimeScope | 800 | 2259ms | 517ms | 2776ms |
+| 1200 | 3205ms | 627ms | 3832ms |
+Efficiency of ToTG-pile Dataset.
+To further analyze the impact of the ToTG-Pile dataset, we retrained the model by excluding it from the training set and reported the corresponding results on short-video temporal benchmark Charades-STA, long-video benchmarks V-STaR and ToTG-bench (long), as shown in Table 9(b).
+It can be observed that the performance gap becomes particularly pronounced on V-STaR and ToTG-bench (long), where ToTG-pile brings performance gains of 40 and 20 points, respectively, compared to the setting without ToTG-pile.
+This demonstrates the crucial value of ToTG-pile in establishing TimeScope’s long-video understanding capability and task-oriented reasoning proficiency.
+| Setting | Max Frames | V-STaR(long) | ToTG-Bench(long) | ||
+| R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | ||
+| w/o Holistic VR. | 800 | 73.2 | 61.4 | 37.8 | 29.5 |
+| w Holistic VR. | 2000 | 80.7 | 76.1 | 41.0 | 37.1 |
+| Method | Charades-STA | V-STaR(long) | ToTG-bench(long) | |||
+| R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
+| W/o ToTG-Pile | 74.9 | 55.8 | 47.7 | 39.7 | 29.4 | 22.7 |
+| W ToTG-Pile | 78.9 | 61.2 | 86.4 | 83.0 | 37.8 | 34.0 |
+6 Qualitative results
+We show the results of TimeScope in Figure 3, which can be seen that TimeScope exhibits robust performance with good coarse-grained segment retrieval and fine-grained temporal grounding capabilities.
+More results can be seen in Appendix.
 7 Conclusion
@@
 We hope this work will stimulate future research on Task-Oriented Temporal Grounding and propel MLLMs toward deeper temporal understanding of video.
-Appendix A Use of LLMs
-In the process of writing this paper, we utilized a Large Language Model (LLM) solely as a tool for polishing the language and enhancing the readability of the manuscript.
-The LLM was employed to refine the wording, grammar, and coherence of the text, ensuring that the content is clear and well-structured.
-However, it is important to note that the LLM did not play any significant role in the conception of research ideas or the formulation of the paper’s content.
-All research ideas, methodologies, and conclusions presented in this paper are the original work of the authors.
-We understand that we bear full responsibility for the contents of this paper, including any text that has been processed by the LLM.
-We have carefully reviewed and verified all content to ensure its accuracy and originality.
-Any contributions made by the LLM were limited to language enhancement and did not involve any form of plagiarism or scientific misconduct.
-Appendix B Benchmark Details
+Appendix A Different and Other Grounding-QA
+Although some traditional groundingQA datasets—such as CG-Bench, NExT-GQA, and ReXTime already require localizing temporal evidence based on the question rather than on explicit event phrases, the Task-oriented Temporal Grounding task we propose is fundamentally different.
+These differences are illustrated in Figure 1.
+Specifically, questions in conventional question-based grounding works usually still contain words strongly correlated with the target time interval, allowing existing temporal-grounding models to localize the interval by relying on the dominant cues in the question.
+In contrast, our task-oriented setting emphasizes reasoning and thinking driven by the question: our questions typically lack descriptors that are directly tied to the answer, and instead demand that the model first infer which temporal segments can solve the question and then ground them.
+Appendix B Holistic Video Representation Detail
+We conducted a comprehensive ablation on the parameters of Holistic Video Representation (HVR), including the chunk size used to partition the video and the compression ratio applied when generating HVR.
+We obtain HVR from Detail Video Representation through a pooling operation.
+During training, we randomly select one-third of the samples to be trained with HVR.
+As shown in Tables 1 and 2, although using HVR leads to a slight performance drop when the video is input at 1 fps with a maximum of 800 frames, we observe clear performance gains when the maximum number of input frames is increased to 2,000.
+With HVR, the model surpasses the original zeroshot results.
+| Method | Max Frames | V-STaR(long) | ToTG-Bench(long) | |||
+| VideoChunk | HVR Rate | R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
+| 0 | 1* | 800 | 73.2 | 61.4 | 37.8 | 29.5 |
+| 0 | 2* | 800 | 63.6 | 63.6 | 31.2 | 27.7 |
+| 0 | 4* | 800 | 62.5 | 60.2 | 30.4 | 26.4 |
+| 150 | 1* | 800 | 61.3 | 60.2 | 29.5 | 26.1 |
+| 150 | 2* | 800 | 63.6 | 63.6 | 31.8 | 27.3 |
+| 150 | 4* | 800 | 62.5 | 61.3 | 30.6 | 26.9 |
+| 300 | 1* | 800 | 60.2 | 58.0 | 28.7 | 24.9 |
+| 300 | 2* | 800 | 63.6 | 63.6 | 32.4 | 26.3 |
+| 300 | 4* | 800 | 61.3 | 59.1 | 28.6 | 25.8 |
+| Method | Max Frames | V-STaR(long) | ToTG-Bench(long) | |||
+| VideoChunk | HVR Rate | R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
+| 0 | 1* | 800 | 73.2 | 61.4 | 37.8 | 29.5 |
+| 0 | 1* | 2000 | - | - | - | - |
+| 150 | 1* | 2000 | 62.5 | 60.2 | 31.4 | 26.7 |
+| 150 | 2* | 2000 | 80.7 | 76.1 | 41.0 | 37.1 |
+| 150 | 4* | 2000 | 86.4 | 81.8 | 43.2 | 39.7 |
+| 300 | 1* | 2000 | 64.7 | 62.5 | 33.4 | 28.2 |
+| 300 | 2* | 2000 | 79.5 | 78.4 | 40.6 | 37.3 |
+| 300 | 4* | 2000 | 81.8 | 80.6 | 42.7 | 38.8 |
+Appendix C ToTG-bench detail
+ToTG-Bench comprises 337 videos and nearly 500 questions.
+As shown in Figure 2, it exhibits rich diversity in task types, video categories, video durations, and the temporal locations of target intervals.
+The average video length is 805 seconds.
+Specifically, the benchmark covers 13 task types—such as action reasoning, OCR perception, and temporal reasoning—spanning a wide range of video-understanding tasks.111We adopt the definitions of task type and video category from Video-MME.
+In addition, it includes 35 video categories from numerous real-world domains.
+Video durations range from a few minutes to one hour, and the target intervals are uniformly distributed along the timeline to ensure an unbiased evaluation.
+Explanations of the 13 Task Categories in ToTG-Bench:
+OR (Object Reasoning): Identify why an object is present and how it causally affects later events in the clip.
+SR (Spatial Reasoning): Infer spatial relationships (e.g., behind, left-of) between objects that are never simultaneously visible.
+TR (Temporal Reasoning): Determine the correct chronological order of events shown out of sequence or with missing segments.
+AR (Action Reasoning): Predict the next action an agent will perform given the current context and goals.
+SP (Spatial Perception): Locate and describe where objects are in the 3-D scene from egocentric or exocentric views.
+AP (Attribute Perception): Recognize object properties such as color, material, or state that remain constant across frames.
+TP (Temporal Perception): Detect when a change or event occurs and report its exact start and end times.
+EG (Ego): Understand what the camera-wearer is doing, attending to, or will do next from first-person video.
+OP (OCR Problems): Read and transcribe text that appears on screens, signs, or documents within the video.
+AT (Action Recognition): Classify which predefined action is being performed in a given segment of the video.
+CP (Counting Problem): Accurately count how many instances of a specified object or person appear throughout the clip.
+TU (Tutorial): Follow instructional video content and answer questions about the steps or techniques demonstrated.
+Appendix D Metric
 To evaluate the performance of our model on temporal grounding and VideoQA tasks, we employ the Intersection-over-Union (IoU) metric and its variants.
@@
 The formal definitions are as follows:
-B.1 IoU and mIoU
+D.1 IoU and mIoU
 The fundamental IoU metric is defined as:
@@
 We also evaluate performance using IoU thresholds, which measure the percentage of predictions exceeding specific IoU values:
-IoU@0.3: Percentage of predictions with IoU 0.3
-IoU@0.5: Percentage of predictions with IoU 0.5
-IoU@0.7: Percentage of predictions with IoU 0.7
-Here, denotes the duration of the overlapping region between the predicted window and the ground truth window .
-represents the duration of their union, while and denote the durations of the predicted and ground truth windows, respectively.
-The indicator function equals 1 when the condition is true and 0 otherwise.
-IoU measures the overall alignment between and , providing a balanced assessment of both precision and recall by considering the overlap relative to their union.
-The threshold-based metrics (IoU@0.3, IoU@0.5, IoU@0.7) evaluate the model’s ability to produce high-quality predictions meeting different precision standards, while mIoU provides an overall average performance measure across all samples.
-B.2 Query Center Robustness
-To measure the effectiveness of balanced query centers in the ToTG-benchmark, we evaluated the sensitivity of state-of-the-art (SOTA) models to query centers.
-As shown in the figure, experiments indicate that most models perform better on test data with query centers positioned earlier rather than later or in the middle.
-We measured the difference between the best and worst performance of various models when the query center varies.
-The results show that Timescope maintains its effectiveness regardless of the position of the query center (for instance, the gap is only 35% at iou@0.3, which is significantly better than Qwen2.5vl’s 78% and 23.8% higher than UniTime).
-This further proves the effectiveness of Timescope’s training.
 | Query Center | 0-0.1 | 0.1-0.2 | 0.2-0.3 | 0.3-0.4 | 0.4-0.5 | 0.5-0.6 | 0.6-0.7 | 0.7-0.8 | 0.8-0.9 | 0.9-1 | diff | |
@@
 | R@0.7 | 0.357 | 0.375 | 0.533 | 0.286 | 0.263 | 0.294 | 0.400 | 0.500 | 0.353 | 0.267 | 50.6% | |
-Appendix C Model Arch and Training detail
-Timescope initializes with the base model of VideoXL2 because it can handle very long sequences, facilitating the construction of long video Task-oriented Temporal Grounding tasks.
-The overall architecture is shown in the figure above.
-The architecture of Timescope consists of four components: a visual encoder, Dynamic Token Synthesis (DTS), an MLP projector, and a Large Language Model (LLM).
-For the visual encoder, we use SigLIP to encode visual inputs into dense visual features.
-The DTS module, located after the visual encoder, is constructed by combining spatial-temporal attention blocks and 3D convolutional layers.
-It aims to process visual features extracted from four consecutive visual inputs as a single group.
-This design has been validated in Video-XL-Pro.
-Next, the architecture applies average pooling to adjacent features to further compress the representation.
-Then, a two-layer MLP projector processes these pooled features, projecting them into the embedding space of the LLM.
-For our LLM, we adopt Qwen2.5-7B.
-Furthermore, in Timescope, to enhance the model’s temporal understanding capability, we prepend explicit timestamp tokens (e.g., Time: 4.0 Second) for every four consecutive frames group across the entire frames sequence.
-This direct timing information substantially improves the model’s temporal awareness.
-During training, we randomly perform shift operations (e.g., 4s to 1004 seconds) and cut operations (e.g., cutting out 10-20 seconds of a video as a single data point) on the explicit timestamp tokens of the input video as a whole.
-This training technique greatly enhances Timescope’s multi-stage understanding capability and robustness to query positions, as shown in Table B.2.
-Appendix D Experimental Settings & Additional Results
-We elaborate on the training and inference details of Timescope.
-The reported hyperparameters cover Stage 1, and 2, as specified in Table 1.
-For the inference details, we emphasize the particular context length for different benchmarks, as shown in Table 2.
 | Hyperparameter | Stage 1 | Stage 2 |
@@
 | V-STaR | 1fps(800 frames) |
-Appendix E VideoQA Task Details
-E.1 VideoQA with Temporal Grounding
+IoU@0.3: Percentage of predictions with IoU 0.3
+IoU@0.5: Percentage of predictions with IoU 0.5
+IoU@0.7: Percentage of predictions with IoU 0.7
+Here, denotes the duration of the overlapping region between the predicted window and the ground truth window .
+represents the duration of their union, while and denote the durations of the predicted and ground truth windows, respectively.
+The indicator function equals 1 when the condition is true and 0 otherwise.
+IoU measures the overall alignment between and , providing a balanced assessment of both precision and recall by considering the overlap relative to their union.
+The threshold-based metrics (IoU@0.3, IoU@0.5, IoU@0.7) evaluate the model’s ability to produce high-quality predictions meeting different precision standards, while mIoU provides an overall average performance measure across all samples.
+D.2 Query Center Robustness
+To measure the effectiveness of balanced query centers in the ToTG-benchmark, we evaluated the sensitivity of state-of-the-art (SOTA) models to query centers.
+As shown in the figure, experiments indicate that most models perform better on test data with query centers positioned earlier rather than later or in the middle.
+We measured the difference between the best and worst performance of various models when the query center varies.
+The results show that Timescope maintains its effectiveness regardless of the position of the query center (for instance, the gap is only 35% at iou@0.3, which is significantly better than Qwen2.5vl’s 78% and 23.8% higher than UniTime).
+This further proves the effectiveness of Timescope’s training.
+Appendix E Experimental Settings & Additional Results
+Appendix F VideoQA Task Details
+F.1 VideoQA with Temporal Grounding
 We use Qwen2-VL-7B as the VideoQA model for answer generation.
@@
 Within each interval, we again uniformly sample 32 frames for answer generation.
-E.2 Prompt template for VideoQA
+F.2 Prompt template for VideoQA
 We use the same prompt template for all multiple-choice VideoQA benchmarks:
 System: You are a helpful assistant. User: <video> Question: <question> Options: (A) <Option_A> (B) <Option_B> (C) <Option_C> (D) <Option_D> Please only give the best option. Best Option: Assistant:
-Appendix F Limitations & Future Work
+Appendix G Limitations & Future Work
 Although Timescope demonstrates exceptional performance on various video temporal grounding and video QA benchmarks, it still has several limitations that warrant further exploration: (i) Timescope is currently constrained to temporal grounding tasks (including traditional temporal grounding tasks and Task-Oriented Temporal Grounding tasks).
@@
 The ultimate objective is to develop MLLMs that seamlessly integrate localization, reasoning, and question-answering into a unified framework.
-Appendix G Broader Impacts Statement
+Appendix H Broader Impacts Statement
 Our research introduces a new problem, called Task-oriented Temporal Grounding (ToTG), to formally conceptualize the aforementioned challenge.
@@
 We believe that ToTG and Timescope can advance the development of video temporal understanding.
-Appendix H Qualitative Results
-In Figure 2–4, we present the qualitative results of Timescope on ToTG-bench and V-STaR.
+Appendix I Qualitative Results
+In Figure 5–7, we present the qualitative results of Timescope on ToTG-bench and V-STaR.
 Timescope demonstrates accurate understanding of questions and the ability to provide temporal grounding.